### v1

# Non-Hermitian expander obtained with Haar distributed unitaries

Abstract
We consider a random quantum channel obtained by taking a selection of $d$ independent and Haar distributed $N$ dimensional unitaries. We follow the argument of Hastings to bound the spectral gap in terms of eigenvalues and adapt it to give an exact estimate of the spectral gap in terms of singular values [1, 2]. This shows that we have constructed a random quantum expander in terms of both singular values and eigenvalues. The lower bound is an analog of the Alon-Boppana bound for $d$-regular graphs. The upper bound is obtained using Schwinger-Dyson equations.
1 Introduction
1.1 Notations and properties of finite dimensional operators
[…]
$|\lambda_{1}(M)|\geq|\lambda_{2}(M)|\geq\cdots\geq|\lambda_{N}(M)|.$
We also denote by $s_{k}(M)$ the $k$-th singular value of the operator $M$, i.e., for all $1\leq k\leq N$ we have $s_{k}(M)=\lambda_{k}(M^{\ast}M)^{1/2}$. In particular, we have:
$\|M\|=s_{1}(M)\geq s_{2}(M)\geq\cdots\geq s_{N}(M).$
In quantum mechanics, we consider the state of a $N$-dimensional system to be a positive semidefinite matrix with trace $1$, usually denoted $\rho$. A transformation of such a state, e.g. due to measurements, is described by a quantum channel $\mathcal{E}:\mathrm{M}_{N}(\operatorname{\mathbb{C}})\rightarrow\mathrm{M}_{N}(\operatorname{\mathbb{C}})$, i.e. linear, completely positive, self-adjoint preserving and trace preserving map [3, 4]. Recall that an operator $\mathcal{T}:\mathrm{M}_{N}(\operatorname{\mathbb{C}})\rightarrow\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ is completely positive if the operator $\mathcal{T}\otimes\operatorname{Id}:\mathrm{M}_{N^{2}}(\operatorname{\mathbb{C}})\rightarrow\mathrm{M}_{N^{2}}(\operatorname{\mathbb{C}})$ is positive. For every self-adjoint preserving and completely positive operator $\mathcal{T}:\mathrm{M}_{N}(\operatorname{\mathbb{C}})\rightarrow\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ one can find $K(1),...,K(d)\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ such that for all $M\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ we have (Choi’s Theorem [3, Theorem 2.21]):
$\mathcal{T}(M)=\sum_{s=1}^{d}K(s)MK(s)^{\ast}.$
The previous (not unique) way of writing the operator $\mathcal{T}$ is called Kraus representation. The smallest integer $d\in\operatorname{\mathbb{N}}$ for which a previous writing is possible for a given self-adjoint preserving and completely positive operator $\mathcal{T}$ is called the Kraus rank of the operator $\mathcal{T}$. Finally, $K(1),...,K(d)$ are called the Kraus operators associated with $\mathcal{T}$. Now one can also consider the quantum channel as a sum of tensor products of matrices. In fact, one can consider the trace defined on $\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ by:
$\tau_{N}:\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ $\rightarrow\operatorname{\mathbb{C}}$
$\mathcal{T}$ $\mapsto\sum_{i,j}\operatorname{Tr}[\mathcal{T}(E_{ij})E_{ji}]=\sum_{i,j}(\mathcal{T}(E_{ij}),E_{ij})$
where $(E_{ij})_{i,j\in[N]}$ is the canonical basis of $\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ considered with the usual scalar product $(A,B)=\operatorname{Tr}(AB^{\ast})$ for $A,B\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$. Then we can set the scalar product on $\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$:
$\langle\cdot,\cdot\rangle:\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))^{2}$ $\rightarrow\operatorname{\mathbb{C}}$
$(\mathcal{T},\mathcal{V})$ $\mapsto\tau_{N}(\mathcal{T}\mathcal{V}^{\ast}).$
To find the correspondence between $\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ and $\mathrm{M}_{N}(\operatorname{\mathbb{C}})\otimes\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ one can consider the maps:
$\mathcal{T}_{A,B}:\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ $\rightarrow\mathrm{M}_{N}(\operatorname{\mathbb{C}})$
$M$ $\mapsto AMB$
for $A,B\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$. We have that the application:
$M_{\cdot}:(\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}})),\langle\cdot,\cdot\rangle)$ $\rightarrow(\mathrm{M}_{N}(\operatorname{\mathbb{C}})\otimes\mathrm{M}_{N}(\operatorname{\mathbb{C}}),(\cdot,\cdot))$
$\mathcal{T}_{A,B}$ $\mapsto A\otimes B^{t}=M_{\mathcal{T}_{A,B}}$
induces an isometry for the norm induced by the respective scalar product. Now for any quantum channel $\mathcal{E}$ one can consider $K(1),...,K(d)$ its Kraus operators for a given Kraus decomposition and we have:
$M_{\mathcal{E}}=\sum_{s=1}^{d}K(s)\otimes\overline{K(s)}.$
The spectral distributions of $\mathcal{E}$ and $M_{\mathcal{E}}$ are then equal.
1.2 Quantum information problematic and previous results
Note that a quantum channel defined in this way preserves the set of quantum states. The eigenvalue of largest modulus $\lambda_{1}(\cdot)$ for a quantum channel is always $1$, and there exists an associated eigenvector which is positive and semidefinite, called fixed state and denoted $\hat{\rho}$ (see [4, Introduction], [5, Chapter 6]). As with Markov chains, the second largest eigenvalue (or second largest singular value) of a quantum channel can be seen as a quantification of the distance between the considered quantum channel $\mathcal{E}$ and the ideal quantum channel that would send any state to the fixed point of the quantum channel $\mathcal{E}_{\hat{\rho}}:\rho\mapsto\operatorname{Tr}[\rho]\hat{\rho}$. In view of these considerations, an interest of quantum mechanics is to construct a quantum channel that has a small second eigenvalue as the dimension grows. Therefore, we introduce here a definition of quantum expanders that depends on the spectral distribution of the channel.
Definition 1.1 (Quantum expander (eigenvalues)).
Let $\epsilon>0$ and $\mathcal{E}=(\mathcal{E}_{N})_{N\in\operatorname{\mathbb{N}}}$ be a sequence of quantum channels such that for all $N$, $\mathcal{E}\in\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$. We say that $\mathcal{E}$ is a $\epsilon$-quantum expander in eigenvalues if there exists $N_{0}\geq 1$ such that for all $N\geq N_{0}$ we have:
$|\lambda_{2}(\mathcal{E})|\leq 1-\epsilon.$
This definition corresponds to the classical graph expander definition when defined by the control of the second largest eigenvalue [6]. One could also consider a control on the second largest singular. We denote by $\Pi_{N}\in\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ the orthogonal projector of rank one on $\mathbb{\operatorname{\mathbb{C}}}\operatorname{Id}$.
Definition 1.2 (Quantum expander (singular values)).
Let $\epsilon>0$ and $\mathcal{E}=(\mathcal{E}_{N})_{N\in\operatorname{\mathbb{N}}}$ be a sequence of quantum channels such that for all $N$, $\mathcal{E}\in\mathcal{B}(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$. We say that $\mathcal{E}$ is a $\epsilon$-quantum expander in singular values if there exists $N_{0}\geq 1$ such that for all $N\geq N_{0}$ we have:
$\|\mathcal{E}-\Pi_{N}\|\leq 1-\epsilon.$
In fact, the previous definition corresponds to the second singular value in the case where the eigenspace of $\lambda_{1}(\mathcal{E})$ is of dimension $1$ and the eigenvector is given by $\hat{\rho}:=\frac{1}{N}\operatorname{Id}$. This is the case if we consider $(\frac{1}{\sqrt{d}}U(s))_{s\in[d]}$ as the Kraus decomposition of $\mathcal{E}$ with $(U(s))_{s\in[d]}$ iid Haar distributed. If we now consider a random quantum channel given by a collection of random matrices $(K(s))_{s\in[d]}$, which will be the Kraus decomposition of the operator, its asymptotic spectral properties can be seen as a free probability result. For example, one could consider that $(K(s))_{s\in[d]}$ are such that $d$ is even and $(K(s))_{s\leq d/2}$ are iid random matrices, setting $K(s+d/2)=K(s)^{\ast}$. In this case, $\mathcal{E}$ is a self-adjoint operator. This case allows many results [6, 1, 2]. Both the Hermitian and non-Hermitian cases are treated in the case where $(K(s))_{s\in[d]}$ are random permutations [7], or Haar distributed unitaries [8]. Indeed, Bordenave and Collins showed:
Theorem 1.3 (Bordenave, Collins 2018,2022).
Let $d\geq 1$ be an integer and $(U(s))_{s\in[d]}\in\mathrm{G}_{N}$ iid Haar distributed where $\mathrm{G}_{N}\in\{\mathrm{U}_{N},\mathrm{O}_{N},\mathrm{S}_{N}\}$. The family $(U(s)\otimes\overline{U}(s)|_{\mathds{1}^{\perp}})_{s\in[d]}$ is strongly asymptotically free and converges to a $d$-family of free Haar unit operators.
Figure 1: Plot of the eigenvalues of $\mathcal{E}$ in the hermitian case for $N=40$ and $d=10$. In red the the the semi-circle law of radius $\lambda_{\mathrm{herm},20}:=2\frac{\sqrt{19}}{20}$.
Recall that the case where $\mathrm{G}_{N}=\mathrm{S}_{N}$ [7] was established before the case where $\mathrm{G}_{N}\in\{\mathrm{U}_{N},\mathrm{O}_{N}\}$ [8]. In both cases, it implies expansion for the random quantum channel $\mathcal{E}$ in both the Hermitian and non-Hermitian cases. To be more precise, in the Hermitian case we have with probability one:
$\lim_{N\rightarrow\infty}\lambda_{2}(\mathcal{E}^{\mathrm{h}})=\frac{2\sqrt{d-1}}{d}$
which gives the expansion in terms of the eigenvalues.In the non-Hermitian case we also have:
$\lim_{N\rightarrow\infty}s_{2}(\mathcal{E})=\frac{2\sqrt{d-1}}{d}$
and thus the expansion with respect to the singular values. Also the previous convergence in the Hermitian case and when $\mathrm{G}_{N}=\mathrm{U}_{N}$ was shown independently by Collins, Gaudreau Lamarre and Male. Their absorption theorem regarding the strong asymptotic freeness of independent Haar distributed matrices [9, Theorem 1.4] implies that $(U(s)\otimes\overline{U}(s))_{s\leq d/2}$ converges in law to a Haar unitary system, and by Lancien, Oliveira Santos and Youssef as a particular case of a larger set of laws that give self-adjoint quantum expanders [6]. The absorption phenomenon for convergence in distribution was proved earlier [10]. Absorption leads to the idea that convergences and convergence rates proved for iid Haar distributed unitary matrices $(U(s))_{s\in[d]}$ may remain true for the representation $(U(s)\otimes U(s))_{s\in[d]}$. For example, the speed in $N^{-k}$ for convergence in expectation for $\mathcal{C}^{4k+7}$ functions applied to polynomials in Haar distributed unitaries [11, Theorem 1.1] proved by Félix Parraud. The absorption result combined with the strong asymptotic freeness [12] implied that the limit object for $\ast$-distribution, spectrum and norm was that given by the Haar free system. Finally, for an explicit universal $c>0$, a convergence speed in $N^{c/d}$ for the norm of polynomials in Haar distributed unitaries is given by [13, Corollary 1.2]. It leads to think that the methods used to prove strong asymptotic freeness in [8] and [7] and the speed given in [13] will give a speed in $N^{-c(d)}$ for the expectation of the norm of the polynomials in $(U(s)\otimes\overline{U}(s))_{s\in[d]}$ and therefore a different convergence speed for $s_{2}(\mathcal{E}^{m})^{1/m}=\|\mathcal{E}^{m}-\Pi_{N}\|^{1/m}$ than the method given here.
1.3 Model, Alon-Boppana bound and main theorem
Throughout this article, for $d=d_{N}\in\mathbb{N}^{\ast}$ sequence of degrees and $(U(s))_{s\in[d]}\in\mathrm{U}_{N}^{d}$ sequence of $d$-tuple unitaries, we consider the following random operator:
$\mathcal{E}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathrm{M}_{N}(\mathbb{C})$ $\rightarrow\mathrm{M}_{N}(\mathbb{C})$
$M$ $\mapsto\frac{1}{d}\sum_{s=1}^{d}U(s)^{\ast}MU(s).$ (1)
This operator is completely positive, unit-preserving, self-adjointness preserving and trace-preserving, and is therefore a quantum channel. We consider the optimal asymptotic second largest eigenvalue in the non-Hermitian case:
$\rho_{d}:=\frac{1}{\sqrt{d}},$
and the optimal asymptotic second largest singular value:
$\sigma_{d}:=\frac{2\sqrt{d-1}}{d}.$
The previous quantum channel verifies analogs of Alon-Boppana bound given in the following lemmas.
Lemma 1.4.
For any $d=d_{N}$ sequence of degree, let $(U(s))_{s\in[d]}\in\mathrm{U}_{N}^{d}$ be any sequence of $d$ unitary matrices and let $\mathcal{E}$ be the operator defined by (1). We set $m=m_{N}:=\lfloor\frac{\ln(N)}{4\ln(d)}\rfloor$. For all $2\leq d\leq N^{1/4}$, for all $N\geq 2$ we have:
$s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}\exp(-\frac{1}{2N}).$
The previous Lemma gives a lower bound for the approximation of $\lambda_{2}(\mathcal{E})$ given by $s_{2}(\mathcal{E}^{m})^{1/m}$ with $N$ fixed. The following Lemma gives the lower bound for $s_{2}(\mathcal{E})$ directly.
Lemma 1.5.
For any $d=d_{N}$ sequence of degree, let $(U(s))_{s\in[d]}\in\mathrm{U}_{N}^{d}$ be any sequence of $d$ unitary matrices and let $\mathcal{E}$ be the operator defined by (1). We set $p=p_{N}=\lfloor\frac{\ln(N)}{2\ln(1/\sigma_{d})}\rfloor$. There exists $c>0$ such that for all sequence of degree $2\leq d=d_{N}\leq N^{1/4}$, for all $N\geq 2$ we have:
$s_{2}(\mathcal{E})\geq\sigma_{d}\left(\frac{c}{\ln(N)^{3/2}}\right)^{1/2p}.$
The previous lemma implies that for all sequence $(d_{N})_{N\geq 2}$ such that:
$\ln(d_{N})<<\frac{\ln(N)}{\ln(\ln(N))}$ (2)
we have that almost surely:
$s_{2}(\mathcal{E})\geq\sigma_{d}\left(1-\epsilon_{N}\right),$
for an explicit $\epsilon_{N}\underset{N\rightarrow\infty}{\rightarrow}0$. The previous lemmas give an upper bound on the best eigenvalue or singular value expansion one can get. The following theorems show that for $(U(s))_{s\in[d]}$ iid Haar distributed unitary matrices of dimension $N$ the operator $\mathcal{E}$ given by (1) is indeed a quantum expander with high probability and its expansion is optimal with respect to the previous lemmas (for eigenvalue and singular value approximation).
Theorem 1.6.
For all $d=d_{N}$ sequence of degrees we consider $(U(s))_{s\in[d]}$ iid Haar distributed unitary matrices of dimension $N$. For all $\epsilon>0$, we have that for $N$ large enough, for all $d=d_{N}$ sequence of degree we have:
$\mathbb{P}\left(|\lambda_{2}(\mathcal{E})|\geq\rho_{d}(1+\epsilon)\right)\leq e^{-\frac{1}{8\sqrt{2}}\ln(1+\epsilon)N^{1/12}},$
where $\mathcal{E}$ is the quantum channel defined by Equation (1).
The previous theorem can be shown directly adapting Hastings proof to the non-Hermitian case [1]. It would then be a direct consequence of the following proposition.
Proposition 1.7.
For all $d=d_{N}$ sequence of degrees we consider $(U(s))_{s\in[d]}$ iid Haar distributed unitary matrices of dimension $N$ and for $N\geq 2$ integer we set $m=m_{N}=\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$. For all $\epsilon>0$, for all sequence of degrees $d=d_{N}$ we have:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)\leq\frac{1}{8}N^{13/6}e^{-\frac{1}{4\sqrt{2}}\ln(1+\epsilon)N^{1/12}},$
where $\mathcal{E}$ is the quantum channel defined by Equation (1). In particular we have that for $N$ large enough:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)\leq e^{-\frac{1}{8\sqrt{2}}\ln(1+\epsilon)N^{1/12}}.$
However Theorem 1.6 can also be seen as a consequence of the following theorem.
Theorem 1.8.
For all $d=d_{N}$ sequence of degrees we consider $(U(s))_{s\in[d]}$ iid Haar distributed unitary matrices of dimension $N$. For all $\epsilon>0$ and for $N$ large enough we have for all sequence of degrees $d_{N}$ and sequence of powers $1\leq m=m_{N}\leq\frac{1}{2}\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)\leq\frac{\sqrt{2}(m+1)}{m}N^{25/12}e^{[2\frac{\ln(m+1)}{m}-\ln(1+\epsilon)]\frac{1}{4\sqrt{2}}N^{1/12}},$
where $\mathcal{E}$ is the quantum channel defined by Equation (1). In particular if furthermore $m_{N}\underset{N\rightarrow\infty}{\rightarrow}\infty$ then for $N$ large enough we have:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)\leq e^{-\frac{1}{8\sqrt{2}}\ln(1+\epsilon)N^{1/12}}.$
Detailing the previous theorem computation in the special case where $m=1$ gives the expansion in terms of singular values.
Theorem 1.9.
For all $d=d_{N}$ sequence of degrees we consider $(U(s))_{s\in[d]}$ iid Haar distributed unitary matrices of dimension $N$. For all $\epsilon>0$ and for $N$ large enough we have for all sequence of degree $d_{N}$:
$\mathbb{P}\left(s_{2}(\mathcal{E})\geq\sigma_{d}(1+\epsilon)\right)\leq N^{25/12}e^{-\ln(1+\epsilon)\frac{1}{4\sqrt{2}}N^{1/12}}.$
where $\mathcal{E}$ is the quantum channel defined by Equation (1).
The advantage of the previous theorems is that, together with Lemma 1.4 and Lemma 1.5, we have the following corollaries, which give convergence either for an approximation of the second largest eigenvalue (i.e., the second largest singular value of the $m$-th power of $\mathcal{E}$) or for the second largest singular value of $\mathcal{E}$.
Corollary 1.10.
Let $\epsilon>0$. For all $2\leq d=d_{N}\leq N^{1/4}$ sequence of degrees we have for $N$ large enough:
$\mathbb{P}\left(\left|s_{2}(\mathcal{E}^{m})^{1/m}-\rho_{d}\right|\geq\rho_{d}\epsilon\right)\leq e^{-\frac{1}{8\sqrt{2}}\ln(1+\epsilon)N^{1/12}}$
where $m=\lfloor 2\frac{\ln(N)}{\ln(d)}\rfloor$.
The upper bound given by Theorem 1.8 gives us the expansion in terms of eigenvalues of $\mathcal{E}$, since we have the following inequality:
$|\lambda_{2}(\mathcal{E})|\leq s_{2}(\mathcal{E}^{m})^{1/m}.$
Also Gelfand Theorem states that:
$\lambda_{2}(\mathcal{E})=\liminf_{m\rightarrow\infty}s_{2}(\mathcal{E}^{m})^{1/m}.$
However, these facts are not sufficient to give direct information about the behavior of $\Delta(\mathcal{E})=|\lambda_{2}(\mathcal{E})-\rho_{d}|$ as the dimension grows. In fact, an upper bound for $\Delta(\mathcal{E})$ would be a consequence of an intermediate result which states that:
$\lambda_{2}(\mathcal{E})\geq\left(1+o(1)\right)s_{2}(\mathcal{E}^{m})^{1/m}$
for some sequence $m=m_{N}$ suitable for Theorem 1.8.
Figure 2: Plot of the eigenvalues of $\mathcal{E}$ for $N=40$ and $d=10$. In red the the circle of radius $\lambda_{10}:=\frac{1}{\sqrt{10}}$.
On another hand we have the following corollary for the convergence of the second largest singular value.
Corollary 1.11.
Let $\epsilon>0$. For all $d=d_{N}$ verifying (2) sequence of degrees we have for $N$ large enough:
$\mathbb{P}\left(\left|s_{2}(\mathcal{E})-\sigma_{d}\right|\geq\sigma_{d}\epsilon\right)\leq e^{-\frac{1}{8\sqrt{2}}\ln(1+\epsilon)N^{1/12}}.$
Figure 3: Plot of the eigenvalues of $|\mathcal{E}|:=\sqrt{\mathcal{E}^{\ast}\mathcal{E}}$ for $N=30$ and $d=10$. In red $x=\bar{\lambda}_{10}=2\frac{\sqrt{d-1}}{d}=\frac{3}{5}$.
Except for Lemma 1.4 and Lemma 1.5, all the proofs will rely on an algorithm based on the Schwinger-Dyson Equations (16) used by Hastings [1], which gives good approximate computations of the expectation of the product of traces of words in Haar distributed unitary matrices (see Section 2.2). The lack of dependence on $d_{N}$ for Proposition 1.7 and Theorems 1.8 and 1.9 is due to the fact that iterations of the Schwinger-Dyson equation give upper bounds on the expectations of traces in monomials of Haar distributed unitaries that depend only on the degree of the original monomial (see Section 2.2, Lemma 3.7).
1.4 Proof of Lemma 1.4 and Lemma 1.5
We start with a proof of the Alon-Boppana bounds given by Lemma 1.4 and Lemma 1.5. Throughout this section we consider that for all sequences of degree $d=d_{N}$, for all $N\geq 1$ we have a collection $(U(s))_{s\in[d]}$ of unitary matrices of dimension $N$. In order to lower bound $s_{2}(\mathcal{E}^{m})^{1/m}$ we set for any dimension $N$ and $m\geq 1$ integers:
$\mathrm{E}_{1}:=\tau_{N}({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})=\sum_{i,j}\langle\mathcal{E}^{m}(E_{ij}),\mathcal{E}^{m}(E_{ij})\rangle=\sum_{a=1}^{N^{2}}s_{a}(\mathcal{E}^{m})^{2},$
where $\mathcal{E}$ is the quantum channel defined by (1). When there is no confusion, for any integer $m\geq 1$ and all $\mathcal{S}=(s_{1},...,s_{2m})\in[d]^{2m}$ we set:
$\mathcal{U}(\mathcal{S})$ $:=U(s_{2m})\cdots U(s_{m+1})U(s_{m})^{\ast}\cdots U(s_{1})^{\ast}=\prod_{j=1}^{m}U(s_{2m-j+1})\prod_{j=m+1}^{2m}U(s_{2m-j+1})^{\ast}$
such that developing the powers of $\mathrm{E}_{1}$ we obtain:
$\mathrm{E}_{1}$ $=\left(\frac{1}{d}\right)^{2m}\sum_{(s_{1},...,s_{2m})\in[d]^{m}}\operatorname{Tr}[U(s_{2m})\cdots U(s_{m+1})U(s_{m})^{\ast}U(s_{m-1})^{\ast}\cdots U(s_{1})^{\ast}]\times$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{Tr}[U(s_{1})U(s_{2})\cdots U(s_{m})U(s_{m+1})^{\ast}\cdots U(s_{2m})^{\ast}]$
$=\left(\frac{1}{d}\right)^{2m}\sum_{(s_{1},...,s_{2m})\in[d]^{2m}}\operatorname{Tr}[\prod_{j=1}^{m}U(s_{2m-j+1})\prod_{j=m+1}^{2m}U(s_{2m-j+1})^{\ast}]\operatorname{Tr}[\prod_{j=1}^{m}U(s_{j})\prod_{j=m+1}^{2m}U(s_{j})^{\ast}]$
$=\sum_{(s_{1},...,s_{2m})\in[d]^{2m}}|\operatorname{Tr}[\mathcal{U}(\mathcal{S})]|^{2}.$
Lemma 1.4 is deterministic and, as for the Alon-Boppana bound in the case of $d$-regular graph, it shows that $1-\rho_{d}$ is the best expansion for eigenvalues (see Definition 1.1) one can get with Kraus operators given by unitary matrices.
Proof of Lemma 1.4.
for any $N\geq 1$ we assume a fixed collection $(U(s))_{s\in[d]}$ of $d$ unitaries of dimension $N$. for any dimension $N$ and integer $m$ using the notations above we have:
$1+N^{2}s_{2}(\mathcal{E}^{m})^{2}$ $\geq 1+(N^{2}-1)\lambda_{2}({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})\geq\mathrm{E}_{1}$
$\geq\frac{1}{d^{2m}}\left(d^{m}N^{2}+\sum_{\begin{subarray}{c}(s_{1},...,s_{m})\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \neq(s_{m+1},...,s_{2m})\end{subarray}}\underbrace{|\operatorname{Tr}[\mathcal{U}(\mathcal{S})]|^{2}}_{\geq 0}\right)\geq\rho_{d}^{2m}N^{2}.$
Therefore we have:
$s_{2}(\mathcal{E}^{m})^{1/m}$ $\geq\rho_{d}\exp\left[{\frac{1}{2m}\ln\left(1-\frac{1}{N^{2}\rho_{d}^{2m}}\right)}\right]$
$\geq\rho_{d}\exp\left[-\frac{1}{2N}\right]$
where for the second line we took $m=\lfloor\frac{\ln(N)}{4\ln(d)}\rfloor$. ∎
In order to lower bound $s_{2}(\mathcal{E})$ we set for any dimension $N$ and $p\geq 1$ integer:
$\mathrm{E}_{3}:=\tau_{N}[({\mathcal{E}^{\ast}}\mathcal{E})^{p}]=\sum_{i,j}\langle({\mathcal{E}^{\ast}}\mathcal{E})^{p}(E_{ij}),E_{ij}\rangle=\sum_{a=1}^{N^{2}}s_{a}(\mathcal{E})^{2p},$
When there is no confusion, for any integer $p\geq 1$ and all $\mathcal{S}=(s^{1},...,s^{2p})\in[d]^{2p}$ we set:
$\mathcal{U}(\mathcal{S})$ $:=U(s^{2p})U(s^{2p-1})^{\ast}\cdots U(s^{2})U(s^{1})^{\ast}$
such that developing the powers of $\mathrm{E}_{3}$ we obtain:
$\mathrm{E}_{3}$ $=\left(\frac{1}{d}\right)^{2p}\sum_{(s^{1},...,s^{2p})\in[d]^{2p}}\operatorname{Tr}[U(s^{2p})U(s^{2p-1})^{\ast}\cdots U(s^{2})U(s_{1})^{\ast}]\operatorname{Tr}[U(s^{1})U(s^{2})^{\ast}\cdots U(s^{2p-1})U(s^{2p})^{\ast}]$
$=\left(\frac{1}{d}\right)^{2p}\sum_{(s_{1},...,s_{2p})\in[d]^{2p}}\operatorname{Tr}[\prod_{t=1}^{p}U(s^{2p-2t+2})^{\ast}U(s^{2p-2t+1})]\operatorname{Tr}[\prod_{t=1}^{p}U(s^{2t-1})^{\ast}U(s^{2t})]$
$=\sum_{(s^{1},...,s^{2p})\in[d]^{2p}}|\operatorname{Tr}[\mathcal{U}(\mathcal{S})]|^{2}.$
Lemma 1.5 is also deterministic and shows that $1-\sigma_{d}$ is the best expansion one can get for singular values (see Definition 1.2).
Proof of Lemma 1.5.
for any $N\geq 1$ we suppose fixed a collection $(U(s))_{s\in[d]}$ of $d$ unitaries of dimension $N$. for any $p\geq 1$ we denote by $N^{\prime}(p,0)$ the number of $(s^{t})_{t\in[2p]}\in[d]^{2p}$ such that
$\mathcal{U}(\mathcal{S})=U(s^{2p})U(s^{2p-1})^{\ast}\cdots U(s^{2})U(s^{1})^{\ast}=\operatorname{Id}.$
Using the notations above we have:
$1+N^{2}s_{2}(\mathcal{E})^{2p}$ $\geq 1+(N^{2}-1)\lambda_{2}(({\mathcal{E}}^{\ast}\mathcal{E})^{p})\geq\mathrm{E}_{3}$
$\geq\frac{1}{d^{2p}}\left(N^{\prime}(0,p)N^{2}+\sum_{(s^{t})_{t\in[2p]}\in[d]^{2p}}\mathds{1}(\mathcal{U}(\mathcal{S})\neq\operatorname{Id})\underbrace{|\operatorname{Tr}[\mathcal{U}(\mathcal{S})]|^{2}}_{\geq 0}\right)\geq\frac{N^{\prime}(p,0)}{d^{2p}}N^{2}.$
We consider the following random walk:
$(s_{t})_{t\in\mathbb{N}}$ $\sim\mathcal{U}([d])$
$X_{0}$ $:=\operatorname{Id}$
$\forall t\in\operatorname{\mathbb{N}},\leavevmode\nobreak\ \forall s\in[d]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(X_{2t+1}=X_{2t}U(s))=\frac{1}{d},\leavevmode\nobreak$ $\leavevmode\nobreak\ \mathbb{P}(X_{2t+2}=X_{2t+1}U(s)^{\ast})=\frac{1}{d}.$
Also we consider the random walk defined as above but with $(U^{\prime}(s))_{s\in[d]}$ iid Haar distributed, that is:
$X_{0}^{\prime}$ $:=\operatorname{Id}$
$\forall t\in\operatorname{\mathbb{N}},\leavevmode\nobreak\ \forall s\in[d]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(X_{2t+1}^{\prime}=X_{2t}^{\prime}U(s)^{\prime})=\frac{1}{d},\leavevmode\nobreak$ $\leavevmode\nobreak\ \mathbb{P}(X_{2t+2}^{\prime}=X_{2t+1}^{\prime}{U(s)^{\prime}}^{\ast})=\frac{1}{d}.$
We denote by $N(p,0)$ the number of $(s^{t})_{t\in[2p]}\in[d]^{2p}$ such that:
$U^{\prime}(s^{2p})U^{\prime}(s^{2p-1})^{\ast}\cdots U^{\prime}(s^{2})U^{\prime}(s_{1})^{\ast}=\operatorname{Id}$
and we have that:
$\mathbb{P}(X_{2p}=\operatorname{Id})=\frac{N^{\prime}(p,0)}{d^{2p}}\geq\mathbb{P}(X_{2p}^{\prime}=\operatorname{Id})=\frac{N(p,0)}{d^{2p}}.$
The random walk $(X_{t}^{\prime})_{t\geq 1}$ is a random walk on a $d$-regular graph. Therefore there exists $C>0$ independent from $d$ such that for any integer $p$ we have:
$N(p,0)\geq C\frac{(2\sqrt{d-1})^{2p}}{p^{3/2}}$
(see [14, Theorem 5.3] and [15, 16]). We now consider:
$p:=\lfloor\frac{\ln(N)}{2\ln(1/\sigma_{d})}\rfloor.$
Then there exists $c^{\prime}>0$ such that for any $N\geq 2$ and $2\leq d\leq N^{1/4}$:
$s_{2}(\mathcal{E})$ $\geq\sigma_{d}\left(\frac{c}{p^{3/2}}-\frac{1}{N^{2}\bar{\lambda}^{2p}}\right)^{1/2p}\geq\sigma_{d}\left(\frac{c}{p^{3/2}}-\frac{1}{N}\right)^{1/2p}$
$\geq\sigma_{d}\left(\frac{c^{\prime}}{\ln(N)^{3/2}}\right)^{1/2p}.$
2 Trace method and Schwinger-Dyson equations
From now on, for all $N\geq 1$, for any sequence $d=d_{N}$ of degree, we consider $(U(s))_{s\in[d]}$ iid Haar distributed unitaries of dimension $N$.
2.1 Strategy
The fact that Theorem 1.6 is a consequence of either Proposition 1.7 or Theorem 1.8 comes from the following inequalities on singular and eigenvalues. for any integer $m\geq 1$ we have:
$|\lambda_{2}(\mathcal{E})|^{2m}=|\lambda_{2}(\mathcal{E}^{m})|^{2}\leq s_{2}(\mathcal{E}^{m})^{2}.$
To prove either Proposition 1.7, Theorem 1.8 or Theorem 1.9 we use Markov inequalities and the trace method. For Proposition 1.7 we will use the fact that for any $\epsilon>0$ and for any integer $m\geq 1$ we have:
$\mathbb{P}\left(|\lambda_{2}(\mathcal{E})|\geq\rho_{d}(1+\epsilon)\right)\leq\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{2}\geq\rho_{d}^{2m}(1+\epsilon)^{2m}\right)\leq\frac{\mathbb{E}(s_{2}(\mathcal{E}^{m})^{2})}{\rho_{d}^{2m}(1+\epsilon)^{2m}}.$
Using the notation of Section 1.4, we bound the right hand side of the previous inequality using the following:
$1+\mathbb{E}(s_{2}(\mathcal{E}^{m})^{2})\leq\sum_{a=1}^{N^{2}}s_{a}(\mathcal{E}^{m})^{2}=\mathbb{E}(\mathrm{E}_{1}).$
Therefore we set for $m$ integer and when there is no confusion:
$\mathbb{E}_{1}:=\mathbb{E}(\mathrm{E}_{1})=\left(\frac{1}{d}\right)^{2m}\sum_{\mathcal{S}=(s_{j})_{j\in[2m]}\in[d]^{2m}}\mathbb{E}_{0}(\mathcal{S}),$ (3)
where for all $\mathcal{S}=(s_{1},...,s_{2m})\in[d]^{2m}$ we denote:
$\mathbb{E}_{0}(\mathcal{S}):=\mathbb{E}(|\operatorname{Tr}[\mathcal{U}(\mathcal{S})]|^{2})=\mathbb{E}(|\operatorname{Tr}[U(s_{2m})\cdots U(s_{m+1})U(s_{m})^{\ast}\cdots U(s_{1})^{\ast}]|^{2}).$ (4)
The Markov inequality above becomes:
$\mathbb{P}\left(|\lambda_{2}(\mathcal{E})|\geq\rho_{d}(1+\epsilon)\right)\leq\frac{\mathbb{E}_{1}-1}{\rho_{d}^{2m}(1+\epsilon)^{2m}}.$ (5)
To prove Theorem 1.8 we will use the fact that for any $m,p\geq 1$ integers, for any $\epsilon>0$ we have:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)=\mathbb{P}\left(\lambda_{2}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p})\geq\rho_{d}^{2mp}(1+\epsilon)^{2mp}\right)\leq\frac{\mathbb{E}\left(\lambda_{2}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p})\right)}{\rho_{d}^{2mp}(1+\epsilon)^{2mp}}.$
In order to bound the right hand side of the previous inequality we set for any $m,p\geq 1$ integers:
$\mathrm{E}_{2}:=\tau_{N}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p})=\sum_{i,j}\langle({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p}(E_{ij}),E_{ij}\rangle=\sum_{a=1}^{N^{2}}\lambda_{a}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m}))^{p}\geq 1+\lambda_{2}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m}))^{p}.$
Developing the powers of $\mathcal{E}$ in the expression of $\mathrm{E}_{2}$ we obtain:
$\mathrm{E}_{2}$ $=\left(\frac{1}{d}\right)^{2mp}\sum_{(s_{j}^{t})_{j\in[m],t\in[2p]}\in[d]^{2mp}}\operatorname{Tr}[\prod_{t=1}^{2p}\prod_{j=1}^{m}U(s_{2m-j+1}^{2p-t+1})^{\epsilon_{2p-t+1}}]\operatorname{Tr}[\prod_{t=1}^{2p}\prod_{j=1}^{m}U(s_{j}^{t})^{-\epsilon_{t}}]$ (6)
where for all $1\leq k\leq p$ we have $\epsilon_{2k}=+$ and $\epsilon_{2k+1}=-$. Therefore for any integers $m,p\geq 1$ and all $\mathcal{S}=(s_{j}^{t})_{t\in[2p],j\in[m]}\in[d]^{2pm}$, when there is no confusion we set:
$\mathcal{U}(\mathcal{S})$ $:=U(s_{m}^{2p})\cdots U(s_{1})U(s_{m}^{2p-1})^{\ast}\cdots U(s_{1}^{2p-1})^{\ast}U(s_{m}^{2p-2})\cdots U(s_{1}^{2})U(s_{m}^{1})^{\ast}\cdots U(s_{1}^{1})^{\ast}$
$=\prod_{t=1}^{2p}\prod_{j=1}^{m}U(s_{2m-j+1}^{2p-t+1})^{\epsilon_{2p-t+1}}$
and $\text{and similarly}\leavevmode\nobreak\ \mathbb{E}_{0}(\mathcal{S})$ $:=\mathbb{E}(\operatorname{Tr}[\mathcal{U}(\mathcal{S})]\operatorname{Tr}[\mathcal{U}(\mathcal{S})^{\ast}]).$ (7)
Taking notations from Equation (6) and (7) we introduce the following expectation:
$\mathbb{E}_{2}$ $:=\mathbb{E}(\mathrm{E}_{2})=\left(\frac{1}{d}\right)^{2mp}\sum_{\mathcal{S}=(s^{t}_{j})_{(t,j)\in[2p]\times[m]}\leavevmode\nobreak\ \in[d]^{2mp}}\mathbb{E}_{0}(\mathcal{S}).$ (8)
The Markov inequality in this case becomes:
$\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}(1+\epsilon)\right)\leq\frac{\mathbb{E}_{2}-1}{\rho_{d}^{2mp}(1+\epsilon)^{2mp}}.$ (9)
Finally in order to prove Theorem 1.9 we consider the expression of $\mathrm{E}_{2}$ given by Equation (6) in the case where $m=1$. Indeed Markov inequality states that for any integer $p\geq 1$:
$\mathbb{P}\left(s_{2}(\mathcal{E})\geq\sigma_{d}(1+\epsilon)\right)=\mathbb{P}\left(\lambda_{2}(({\mathcal{E}^{\ast}}\mathcal{E})^{p})\geq\sigma_{d}^{2p}(1+\epsilon)^{2p}\right)\leq\frac{\mathbb{E}\left(\lambda_{2}(({\mathcal{E}^{\ast}}\mathcal{E})^{p})\right)}{\sigma_{d}^{2p}(1+\epsilon)^{2p}}.$
Also for any $p\geq 1$ we have on one hand:
$\mathrm{E}_{3}:=\tau_{N}(({\mathcal{E}^{\ast}}\mathcal{E})^{p})=\sum_{i,j}\langle({\mathcal{E}^{\ast}}\mathcal{E})^{p}(E_{ij}),E_{ij}\rangle=\sum_{a=1}^{N^{2}}\lambda_{a}({\mathcal{E}^{\ast}}\mathcal{E})^{p}\geq 1+\lambda_{2}({\mathcal{E}^{\ast}}\mathcal{E})^{p}.$
On the other hand considering Equation (6) in the case $m=1$ we have:
$\mathrm{E}_{3}$ $=\left(\frac{1}{d}\right)^{2p}\sum_{(s^{t})_{t\in[2p]}\in[d]^{2p}}\operatorname{Tr}[\prod_{t=1}^{2p}U(s^{2p-t+1})^{\epsilon_{2p-t+1}}]\operatorname{Tr}[\prod_{t=1}^{2p}U(s^{t})^{-\epsilon_{t}}]$ (10)
where for any $1\leq k\leq p$ we have $\epsilon_{2k}=+$ and $\epsilon_{2k+1}=-$. Therefore for any integer $p\geq 1$ and all $\mathcal{S}=(s^{t})_{t\in[2p]}\in[d]^{2p}$, when there is no confusion, we set:
$\mathcal{U}(\mathcal{S})$ $:=U(s^{2p})U(s^{2p-1})^{\ast}\cdots U(s^{2})U(s^{1})^{\ast}$
and $\text{and similarly}\leavevmode\nobreak\ \mathbb{E}_{0}(\mathcal{S})$ $:=\mathbb{E}(\operatorname{Tr}[\mathcal{U}(\mathcal{S})]\operatorname{Tr}[\mathcal{U}(\mathcal{S})^{\ast}]).$ (11)
Taking notations from Equation (10) and (11) we introduce the following expectation:
$\mathbb{E}_{3}$ $:=\mathbb{E}(\mathrm{E}_{3})=\left(\frac{1}{d}\right)^{2p}\sum_{\mathcal{S}=(s^{t})_{t\in[2p]}\leavevmode\nobreak\ \in[d]^{2p}}\mathbb{E}_{0}(\mathcal{S}).$ (12)
In this case Markov inequality becomes:
$\mathbb{P}\left(s_{2}(\mathcal{E})\geq\sigma_{d}(1+\epsilon)\right)\leq\frac{\mathbb{E}_{3}-1}{\sigma_{d}^{2p}(1+\epsilon)^{2p}}.$ (13)
We then apply the Hastings strategy, i.e. we iterate the Schwinger-Dyson equation (see Section 2.2) over the expressions (4), (7) and (11), in order to upper bound $\mathbb{E}_{1}$, $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$ (respectively given by Equations (3),(8) and (12)). The expectations defined by Equations (3),(8) and (12) make the notation $\mathbb{E}_{0}(\mathcal{S})$ inconsistent since we have given the same notation for three different definitions. This is only for notational convenience, and since we will be upper-bounding each $\mathbb{E}_{i}$ in independent Sections (Section 5, Section 6 and 7 respectively).
Let us now assume that Theorems 1.8 and 1.9 are proved, and explain why Corollaries 1.10 and 1.11 hold. The probability in Corollary 1.10 verifies:
$\mathbb{P}\left(|s_{2}(\mathcal{E}^{m})^{1/m}-\rho_{d}|>\epsilon\right)=\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\geq\rho_{d}+\epsilon\right)+\mathbb{P}\left(s_{2}(\mathcal{E}^{m})^{1/m}\leq\rho_{d}-\epsilon\right).$
Under the hypothesis on $d=d_{N}$, $m=m_{N}$ and $(U(s))_{s\in[d]}$ of Corollary 1.10 the first probability on the right side of the inequality is bounded by applying Theorem 1.8 and the second is equal to zero for $N$ large enough. Likewise the probability in Corollary 1.11 verifies:
$\mathbb{P}\left(|s_{2}(\mathcal{E})-\sigma_{d}|>\epsilon\right)=\mathbb{P}\left(s_{2}(\mathcal{E})\geq\sigma_{d}+\epsilon\right)+\mathbb{P}\left(s_{2}(\mathcal{E})\leq\sigma_{d}-\epsilon\right).$
Under the hypothesis on $d=d_{N}$ and $(U(s))_{s\in[d]}$ of Corollary 1.11 the first probability on the right side of the inequality is bounded by applying Theorem 1.9 and the second is equal to zero for $N$ large enough.
2.2 Schwinger-Dyson Equations
In this section we consider $k\geq 1$ and $m_{1},...,m_{k}$ integers. For all $1\leq\ell\leq k$ we consider $\mathcal{S}(\ell)=((s_{\ell 1},\epsilon_{\ell 1}),...,(s_{\ell m_{\ell}},\epsilon_{\ell m_{\ell}}))$ a sequence of $([d]\times\{+,-\})^{m_{\ell}}$ and we set:
$\mathcal{U}({\mathcal{S}(\ell)}):=U(s_{\ell 1})^{\epsilon_{\ell 1}}U(s_{\ell 2})^{\epsilon_{\ell 2}}\cdots U(s_{\ell{m_{\ell}}})^{\epsilon_{\ell m_{\ell}}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ L_{\ell}:=\operatorname{Tr}[\mathcal{U}(\mathcal{S})].$ (14)
Denoting $\mathcal{S}=(\mathcal{S}(1),...,\mathcal{S}(\ell))$ we set:
$\mathbb{E}^{\prime}(\mathcal{S}):=\mathbb{E}(L_{1}\cdots L_{k}).$ (15)
Remark 2.1.
One can notice that the definition of $\mathbb{E}^{\prime}$ given in (15) is linked with the definition of $\mathbb{E}_{0}$ used in the previous section given by (7) and (4). Let $\mathcal{S}=(s_{j}^{t})_{j\in[m],t\in[2p]}$ be a sequence of $[d]^{2mp}$. Then one has to consider $\tilde{\mathcal{S}}=(\tilde{\mathcal{S}}(1),\tilde{\mathcal{S}}(2))$ defined by:
$\tilde{\mathcal{S}}(1)$ $=((s_{m}^{2p},+),...,(s_{1}^{2p},+),(s_{m}^{2p-1},-),...,(s_{j+1}^{t},\epsilon_{t}),(s_{j}^{t},\epsilon_{t}),...,(s_{1}^{1},-))$
$\tilde{\mathcal{S}}(2)$ $=((s_{1}^{1},+),(s_{2}^{1},+),...,(s_{m}^{1},+),(s_{1}^{2},-),...,(s_{j}^{t},-\epsilon_{t}),(s_{j+1}^{t},-\epsilon_{t}),...,(s_{m}^{2p},-))$
where we set $\epsilon_{t}=+$ for $t$ even and $\epsilon_{t}=-$ for $t$ odd. We then have that $\mathcal{U}(\tilde{\mathcal{S}}(1))=\mathcal{U}(\tilde{\mathcal{S}}(2))^{\ast}$ and:
$\mathbb{E}_{0}(\mathcal{S})=\mathbb{E}^{\prime}(\tilde{\mathcal{S}}).$
The following proposition states the Schwinger-Dyson equation given by Hastings [1, Equation (19,20)].
Proposition 2.1 (Schwinger-Dyson equation).
Let $(U(s))_{s\in[d]}$ iid Haar distributed $N$-dimensional unitary matrices. Let $k$ and $m_{1},...,m_{k}$ be integers. For all $1\leq\ell\leq k$ let $\mathcal{S}(\ell)=(s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]}\in([d]\times\{+,-\})^{m_{\ell}}$. We set for all $2\leq j\leq m_{1}$:
$L_{1}^{1}(1,j,+)$ $:=\operatorname{Tr}[U(s_{11})^{\epsilon_{11}}\cdots U(s_{1j-1})^{\epsilon_{1j-1}}]$
$L_{1}^{2}(1,j,+)$ $=\operatorname{Tr}[U(s_{1j})^{\epsilon_{1j}}\cdots U(s_{1m_{1}})^{\epsilon_{1m_{1}}}]$
$L_{1}^{1}(1,j,-)$ $:=\operatorname{Tr}[U(s_{11})^{\epsilon_{11}}\cdots U(s_{1j})^{\epsilon_{1j}}]$
$L_{1}^{2}(1,j,-)$ $:=\operatorname{Tr}[U(s_{1j+1})^{\epsilon_{1j+1}}\cdots U(s_{1m_{1}})^{\epsilon_{1m_{1}}}]$
and for all $1\leq\ell\leq k$ and all $1\leq j\leq m_{\ell}$:
$L_{1}(\ell,j,+)$ $:=\operatorname{Tr}[U(s_{11})^{\epsilon_{11}}\cdots U(s_{1m_{1}})^{\epsilon_{1m_{1}}}U(s_{\ell j})^{\epsilon_{\ell j}}\cdots U(s_{\ell m_{\ell}})^{\epsilon_{\ell m_{\ell}}}U(s_{\ell 1})^{\epsilon_{\ell 1}}\cdots U(s_{\ell j-1})^{\epsilon_{\ell j-1}}]$
$L_{1}(\ell,j,-)$ $=\operatorname{Tr}[U(s_{11})^{\epsilon_{11}}\cdots U(s_{1m_{1}})^{\epsilon_{1m_{1}}}U(s_{\ell j+1})^{\epsilon_{\ell j+1}}\cdots U(s_{\ell m_{\ell}})^{\epsilon_{\ell m_{\ell}}}U(s_{\ell 1})^{\epsilon_{\ell 1}}\cdots U(s_{\ell j})^{\epsilon_{\ell j}}]$
Using notations of Equations (14) and (15) we have the following equality:
$\mathbb{E}^{\prime}(\mathcal{S})$ $=\mathbb{E}(\operatorname{Tr}[U(s_{11})^{\epsilon_{11}}U(s_{12})^{\epsilon_{12}}\cdots U(s_{1{m_{1}}})^{\epsilon_{1m_{1}}}]L_{2}\cdots L_{k})$ (16)
$=-\frac{1}{N}\sum_{j=2}^{m_{1}}\mathds{1}((s_{11},\epsilon_{11})=(s_{1j},\epsilon_{1j}))\mathbb{E}(L_{1}^{1}(1,j,+)L_{1}^{2}(1,j,+)L_{2}\cdots L_{k})$ (17)
$+\frac{1}{N}\sum_{j=2}^{m_{1}}\mathds{1}((s_{11},\epsilon_{11})=(s_{1j},-\epsilon_{1j}))\mathbb{E}(L_{1}^{1}(1,j,-)L_{1}^{2}(1,j,-)L_{2}\cdots L_{k})$ (18)
$-\frac{1}{N}\sum_{\ell=2}^{m_{\ell}}\sum_{j=1}^{m_{\ell}}\mathds{1}((s_{11},\epsilon_{11})=(s_{\ell j},\epsilon_{\ell j}))\mathbb{E}(L_{1}(\ell,j,+)L_{2}\cdots L_{\ell-1}L_{\ell+1}L_{k})$ (19)
$=+\frac{1}{N}\sum_{\ell=2}^{m_{\ell}}\sum_{j=1}^{m_{\ell}}\mathds{1}((s_{11},\epsilon_{11})=(s_{\ell j},-\epsilon_{\ell j}))\mathbb{E}(L_{1}(\ell,j,-)L_{2}\cdots L_{\ell-1}L_{\ell+1}L_{k}).$ (20)
In order to prove the previous equations one needs to consider the matrices:
$\Delta_{lk}=\frac{1}{\sqrt{2}}(E_{lk}+E_{kl}),\leavevmode\nobreak\ \leavevmode\nobreak\ \delta_{lk}=\frac{i}{\sqrt{2}}(E_{lk}-E_{kl}).$
The set of these matrices for all $(l,k)\in[N]^{2}$ is a basis for the $\mathbb{R}$- vector space of $N$-dimensional hermitian matrices. We denote by $\Theta$ this basis.
Lemma 2.2.
Let $X,Y\in\mathrm{M}_{N}(\mathbb{C})$. We have:
$\sum_{T\in\Theta}\operatorname{Tr}(XT^{2})=N\operatorname{Tr}(X)$
$\sum_{T\in\Theta}\operatorname{Tr}(XTYT)=\operatorname{Tr}(X)\operatorname{Tr}(Y)$
$\sum_{T\in\Theta}\operatorname{Tr}(XT)\operatorname{Tr}(YT)=\operatorname{Tr}(XY).$
Proof of Schwinger-Dyson Equation 2.1.
Let $s_{0}\in[d]$ such that $s_{0}=s_{11}$ and any $T\in\Theta$. For $t\geq 0$, we operate the change of variable:
$U(s_{0})\rightarrow\mathrm{e}^{itT}U(s_{0})^{\epsilon_{11}}$
in the expectation $\mathbb{E}(\operatorname{Tr}[X_{11}X_{12}\cdots X_{1m_{1}}]L_{2}\cdots L_{k})$. For all $(\ell,j)$ we set:
$\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (s_{\ell j},\epsilon_{\ell j})=(s_{11},\epsilon_{11}):\leavevmode\nobreak\ \leavevmode\nobreak\ X_{\ell j}^{t}=\mathrm{e}^{itT}U(s_{0})^{\epsilon_{11}}$
$\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ (s_{\ell j},\epsilon_{\ell j})=(s_{11},-\epsilon_{11}):\leavevmode\nobreak\ \leavevmode\nobreak\ X_{\ell j}^{t}=U(s_{0})^{-\epsilon_{11}}\mathrm{e}^{-itT}$
$\text{otherwise}:\leavevmode\nobreak\ \leavevmode\nobreak\ X_{\ell j}^{t}=X_{\ell j}=U(s_{\ell j})^{\epsilon_{\ell j}}.$
For all $t\geq 0$ and $1\leq\ell\leq k$ we set $L_{\ell}^{t}:=\operatorname{Tr}[X_{\ell 1}^{t}\cdots X_{\ell m_{\ell}}^{t}]$. Due to the invariance of the Haar measure under unitary deterministic transformation we have $\operatorname{Tr}[TX_{\ell 1}^{t}\cdots X_{\ell m_{\ell}}^{t}]L_{2}^{t}\cdots L_{k}^{t}\sim\operatorname{Tr}[TX_{\ell 1}\cdots X_{\ell m_{\ell}}]L_{2}\cdots L_{k}$. Therefore, the first order expression in $t$ of $\mathbb{E}(\operatorname{Tr}[TX_{\ell 1}^{t}\cdots X_{\ell m_{\ell}}^{t}]L_{2}^{t}\cdots L_{k}^{t})$ is:
$0=$ $\mathbb{E}(\operatorname{Tr}[T^{2}X_{11}\cdots X_{1m_{1}}]L_{2}\cdots L_{k})$
$+\sum_{j=2}^{m_{1}}\mathds{1}((s_{11},\epsilon_{11})=(s_{1j},\epsilon_{1j}))\mathbb{E}(\operatorname{Tr}[TX_{11}\cdots X_{1j-1}TX_{1j}X_{1j+1}\cdots X_{1m_{1}}]L_{2}\cdots L_{k}]$
$-\sum_{j=2}^{m_{1}}\mathds{1}((s_{11},\epsilon_{11})=(s_{1j},-\epsilon_{1j}))\mathbb{E}(\operatorname{Tr}[TX_{11}\cdots X_{1j}TX_{1j+1}\cdots X_{1m_{1}}]L_{2}\cdots L_{k}]$
$+\sum_{\ell=2}^{k}\sum_{j=1}^{m_{\ell}}\mathds{1}((s_{11},\epsilon_{11})=(s_{\ell j},\epsilon_{\ell j}))\mathbb{E}(\operatorname{Tr}[TX_{11}\cdots X_{1m_{1}}]\operatorname{Tr}[TX_{\ell 1}\cdots X_{\ell j-1}TX_{\ell j}\cdots X_{\ell m_{\ell}}]L_{2}\cdots L_{\ell-1}L_{\ell+1}\cdots L_{k}]$
$-\sum_{\ell=2}^{k}\sum_{j=1}^{m_{1}}\mathds{1}((s_{11},\epsilon_{11})=(s_{1j},-\epsilon_{\ell j}))\mathbb{E}(\operatorname{Tr}[TX_{11}\cdots X_{1m_{1}}]\operatorname{Tr}[X_{\ell 1}\cdots X_{\ell j}TX_{1j+1}\cdots X_{\ell m_{\ell}}]L_{2}\cdots L_{\ell-1}L_{\ell+1}\cdots L_{k}].$
Applying Lemma 2.2, we obtain the desired equations. ∎
3 Coding iterations of Schwinger Dyson equations
Applying the Schwinger-Dyson equation to any term of the form $\mathbb{E}(L_{1}\cdots L_{k})$ of Section 2.2 gives rise to terms $\mathbb{E}(L_{1}^{\prime}\cdots L_{k^{\prime}}^{\prime})$ of the same form, with the same original matrices $(s_{\ell j},\epsilon_{\ell j})$ up to some cancellations and permutations of their position in the expressions of $L^{\prime}_{\ell^{\prime}}$. This can be seen as a random process obtained by an algorithm. Before giving the algorithm, we introduce some definitions and notations.
Definition 3.1.
Let $p,p^{\prime}\geq 0$ be integers. We call $\mathcal{S}=((s_{1},\epsilon_{1}),...,(s_{p},\epsilon_{p}))\in([d]\times\{+,-\})^{p}$ and $\mathcal{S}^{\prime}=((s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{p^{\prime}}^{\prime},\epsilon_{p^{\prime}}^{\prime}))\in([d]\times\{+,-\})^{p^{\prime}}$ two $1$-words. We call $p$ (resp $p^{\prime}$) the length of the $1$-word $\mathcal{S}$ (resp. $\mathcal{S}^{\prime}$). We say that these two words are equivalent if there exists $o\in\mathbb{N}$ such that:
$U(s_{o+1})^{\epsilon_{o+1}}U(s_{o+2})^{\epsilon_{o+2}}\cdots U(s_{p})^{\epsilon_{p}}U(s_{1})^{\epsilon_{1}}\cdots U(s_{o})^{\epsilon_{o}}=U(s_{1}^{\prime})^{\epsilon_{1}^{\prime}}\cdots U(s_{m^{\prime}}^{\prime})^{\epsilon_{m^{\prime}}^{\prime}}.$
For all $1$-word $\mathcal{S}$ we denote by $\ell(\mathcal{S})$ the minimal length for a $1$-word $\mathcal{S}^{\prime}$ equivalent to $\mathcal{S}$. We then call $\mathcal{S}^{\prime}=((s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{p^{\prime}}^{\prime},\epsilon_{p^{\prime}}^{\prime}))\sim\mathcal{S}$ with length $p^{\prime}=\ell(\mathcal{S})$ a minimal writing of the equivalent class of $\mathcal{S}$. For a given minimal writing $\mathcal{S}^{\prime}$, all minimal writings are obtained by considering all $((s_{o+1}^{\prime},\epsilon_{o+1}^{\prime}),...,(s_{p^{\prime}}^{\prime},\epsilon_{p^{\prime}}^{\prime}),(s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{o}^{\prime},\epsilon_{o}^{\prime}))$. We say that $\mathcal{S}^{\prime}$ is the minimal writing of $\mathcal{S}$ if $((s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{p^{\prime}}^{\prime},\epsilon_{p^{\prime}}^{\prime}))$ is an non-decreasing sub-sequence of $((s_{1},\epsilon_{1}),...,(s_{p},\epsilon_{p}))$. Finally we say that the $1$-word $\mathcal{S}$ is of trivial trace if it is equivalent to the empty word, i.e. when for some integer $o\in\mathbb{N}$ we have that:
$U(s_{o+1})^{\epsilon_{o+1}}U(s_{o+2})^{\epsilon_{o+2}}\cdots U(s_{p})^{\epsilon_{p}}U(s_{1})^{\epsilon_{1}}\cdots U(s_{o})^{\epsilon_{o}}=\mathrm{Id}.$
In general we set:
$\mathcal{U}(\mathcal{S}):=U(s_{1})^{\epsilon_{1}}\cdots U(s_{p})^{\epsilon_{p}}.$
We call $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\leq k}$ and $\mathcal{S}^{\prime}=((s_{\ell j}^{\prime},\epsilon_{\ell j}^{\prime})_{j\in[m_{\ell}^{\prime}]})_{\ell\leq k}$ two $k$-words. We say that these two $k$-words are equivalent if for all $1\leq\ell\leq k$ we have $\mathcal{S}(\ell)=((s_{\ell 1},\epsilon_{\ell 1}),...,(s_{\ell m_{\ell}},\epsilon_{\ell m_{\ell}}))\sim\mathcal{S}^{\prime}(\ell)=((s_{\ell 1}^{\prime},\epsilon_{\ell 1}^{\prime}),...,(s_{\ell m_{\ell}^{\prime}}^{\prime},\epsilon_{\ell m_{\ell}^{\prime}}^{\prime}))$. We say that $\mathcal{S}^{\prime}$ is a (resp. the) minimal writing of the equivalent class if for all $1\leq\ell\leq k$, $\mathcal{S}^{\prime}(\ell)$ is a (resp. the) minimal writing of its equivalence class. All minimal writings are called minimal word. Finally we say that the $k$-word $\mathcal{S}$ is of trivial trace if for all $1\leq\ell\leq k$, $\mathcal{S}(\ell)$ is of trivial trace.
Remark 3.1.
These notions take their interest from the fact that, when iterating the Schwinger-Dyson equations, we can consider that we are applying the equation to some minimal $k$-word $\mathcal{S}$ which gives rise to other $k^{\prime}$-words $\mathcal{S}^{\prime}$. Before reapplying the Schwinger-Dyson equation, one must choose one of the resulting $\mathcal{S}^{\prime}$ and consider its minimal writing in order to proceed. In particular it is important to notice that for $\mathcal{S}=((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))$ and $\mathcal{S}^{\prime}=((s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{m^{\prime}}^{\prime},\epsilon_{m^{\prime}}^{\prime}))$ two equivalent $1$-words, we do not necessarily have $\mathcal{U}(\mathcal{S})=\mathcal{U}(\mathcal{S}^{\prime})$ but:
$\operatorname{Tr}(\mathcal{U}(\mathcal{S}))=\operatorname{Tr}(\mathcal{U}(\mathcal{S}^{\prime})).$
Example 3.1.
Let $U(1),U(2),U(3)$ be three independent Haar distributed unitaries. We set $\mathcal{S}=((1,+),(1,-),(2,+),(3,-),(2,+),(2,-))$. The easiest way to proceed is to consider:
$\mathcal{U}(\mathcal{S}):=U(1)U(1)^{\ast}U(2)U(3)^{\ast}U(2)U(2)^{\ast}.$
  • The word $\mathcal{S}$ is equivalent to $\mathcal{S^{\prime}}=((2,+),(3,-),(2,+),(2,-))$ (we have $\mathcal{U}(\mathcal{S}^{\prime}):=U(2)U(3)^{\ast}U(2)U(2)^{\ast}$)
  • A minimal writing of $\mathcal{S}$ is $\mathcal{S^{\prime}}=((2,+),(3,-))$ (we have $\mathcal{U}(\mathcal{S}^{\prime}):=U(2)U(3)^{\ast}\neq\mathcal{U}(\mathcal{S})$)
  • the minimal writing of $\mathcal{S}$ is $\mathcal{S^{\prime}}=((3,-),(2,+))$.
3.1 Encoding the matrices movements
Throughout this section we fix a minimal $k$-word $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\leq k}$, where $k$ and $m_{1},...,m_{k}$ are fixed integers (see Definition 3.1). We denote:
$m:=\sum_{\ell=1}^{k}m_{\ell}$ (21)
The goal of this section is to encode the movements of the matrix $(s_{\ell j},\epsilon_{\ell j})\sim U(s_{\ell j})^{\epsilon_{\ell j}}$ originally at position $(\ell,j)$ after successive iterations of the Schwinger-Dyson equations applied to (15). We initiate our algorithm by setting the original path:
$e_{0}=((1,j)_{j\in[m_{1}]},...,(k,j)_{j\in[m_{k}]})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ p_{0}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon^{(0)}=1$
Also we set:
$\overline{e}_{0}=\mathbb{E}^{\prime}(\mathcal{S})$
where $\mathbb{E}^{\prime}(\cdot)$ is define by (15). The idea is to give an explicit algorithm that gives us the evolution of the traces on the right side of the equation when we iterate Schwinger-Dyson equations. For all $n\geq 0$ we define for all sequences $(\ell_{1},\ell_{2},...,\ell_{n})\in\mathbb{N}^{n}$, $(j_{1},...,j_{n})\in\mathbb{N}^{n}$ and $(\epsilon_{1},...,\epsilon_{n})\in\{+,-\}^{n}$ the sequence of paths $e((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})$ following the coming algorithm. We initiate our algorithm with $n=1$ and a triplet $\mathrm{P}=(\ell_{1},j_{1},\epsilon_{1})$ that we will call a pattern of length $1$.
If $\ell_{1}=1$ and $j_{1}=1$, or $\ell_{1}\geq k+1$, or $j_{1}\geq m_{\ell_{1}}+1$ or $s_{11}\neq s_{\ell_{1},j_{1}}$ we set:
$e(\ell_{1},j_{1},\epsilon_{1})=\emptyset\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{e}(\ell_{1},j_{1},\epsilon_{1})=0.$
If $\ell_{1}=1$, $1\leq j_{1}\leq m_{1}$ and $s_{11}=s_{\ell_{1},j_{1}}$ then:
  • if we have $\epsilon_{1}=+$ and $\epsilon_{11}=\epsilon_{1}\epsilon_{1j_{1}}=\epsilon_{1j_{1}}$, i.e. when we consider a term in the first line of Schwinger-Dyson equation (17), then we set:
    $\epsilon^{(1)}$ $=-1$
    $e(1,j_{1},\epsilon_{1})$ $=((1,j)_{j\leq{j_{1}-1}},(1,j)_{j_{1}\leq j\leq m_{1}},(2,j)_{j\in[m_{2}]},...,(k,j)_{j\in[m_{k}]}),$
    $e_{rs}(1,j_{1},\epsilon_{1})$ $=((1,j)_{j\leq{j_{1}-1}},(2,j)_{j\leq m_{1}-j_{1}+1},(3,j)_{j\in[m_{2}]},...,(k+1,j)_{j\in[m_{k}]})$
    $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})$ $=((s_{1j},\epsilon_{1j})_{j\leq j_{1}-1},(s_{1j},\epsilon_{1j})_{j_{1}\leq j\leq m_{1}},(s_{2j},\epsilon_{2j})_{j\leq m_{2}},...,(s_{kj},\epsilon_{kj})_{j\in[m_{k}]}).$ (22)
    We call $e(\ell_{1},j_{1},\epsilon_{1})$ the path of the pattern $(\ell_{1},j_{1},\epsilon_{1})$, $e_{rs}$ its rescaled path and $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})$ the word of generation 1. In this case and with respect to $\mathcal{S}$ minimality, for all $11\leq\ell\leq k+$ 1 the word $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})(\ell)$ can not be of trivial trace. Therefore in this case we set:
    $p_{1}=p_{0}=0.$
    We also define the function:
    $f_{1}:$ $(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$ $\rightarrow(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$
    $(1,j)$ $\mapsto(1,j)\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq j_{1}-1$
    $(1,j)$ $\mapsto(2,j-j_{1})\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j_{1}\leq j\leq m_{1}$
    $(\ell,j)$ $\mapsto(\ell+1,j)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ \leavevmode\nobreak\ 2\leq\ell\leq k\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and all}\leavevmode\nobreak\ \leavevmode\nobreak\ 1\leq j\leq m_{\ell}.$
    We call $f_{1}$ the first tracking function.
    We denote by $\mathcal{S}_{min}(\cdot)$ the minimal writing of $\mathcal{S}()$, that is:
    $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})=((s_{1j}^{\prime},\epsilon_{1j}^{\prime})_{j\leq m_{1}^{\prime}},...,(s_{k+1j}^{\prime},\epsilon_{k+1j}^{\prime})_{j\leq m_{k+1}^{\prime}})$
    where we have for all $11\leq\ell\leq k+$ 1, $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})(\ell)$ is the minimal writing for $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})(\ell)$ defined by Equation (22). The integers $m_{1}^{\prime},...,m_{k+1}^{\prime}$ are completely determined the simplifications in $\mathcal{U}(\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})(\ell))$, i.e. when in the product given by (14) we see $U(s)^{\epsilon}U(s)^{-\epsilon}$. We set furthermore the minimal path for $(\ell_{1},j_{1},\epsilon_{1})$ as:
    $e_{min}(\ell_{1},j_{1},\epsilon_{1})=((1,j)_{j\leq m_{1}^{\prime}},(2,j)_{j\leq m_{2}^{\prime}},(3,j)_{j\leq m_{3}^{\prime}}).$
    Finally we define the mean of the path $(\ell_{1},j_{1},\epsilon_{1})$:
    $\bar{e}(\ell_{1},j_{1},\epsilon_{1})$ $=\epsilon^{(1)}\frac{N^{p_{1}}}{N}\mathbb{E}^{\prime}(\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1}))$
    where we recall that $\mathbb{E}^{\prime}(\cdot)$ is defined by Equation (15). By definition of the minimal writing, there exists strictly non-decreasing functions $\Phi_{1}:[m_{1}^{\prime}]\rightarrow[j_{1}-1]$ and $\Phi_{2}:[m_{2}^{\prime}]\rightarrow[j_{1};m_{1}]$ such that for all $1\leq j\leq m_{1}^{\prime}$, $(s^{\prime}_{1j},\epsilon_{1j}^{\prime})=(s_{1\Phi_{1}(j)},\epsilon_{1\Phi_{1}(j)})$ and likewise for $\ell=2$. We then define the rescaled tracking function:
    $f^{rs}_{1}:$ $(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$ $\rightarrow(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$
    $(1,j)$ $\mapsto(1,\Phi^{-1}_{1}(j))\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ j\in\Phi_{1}([m_{1}^{\prime}])$
    $(1,j)$ $\mapsto(2,\Phi^{-1}_{2}(j))\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ j\in\Phi_{2}([m_{2}^{\prime}])$
    $(\ell,j)$ $\mapsto(\ell+1,j)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ \leavevmode\nobreak\ 2\leq\ell\leq k+1\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and all}\leavevmode\nobreak\ \leavevmode\nobreak\ 1\leq j\leq m_{\ell}$
    $(\ell,j)$ $\mapsto(0,0)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{otherwise}.$
    If $\epsilon_{1}=+$ and $\epsilon_{11}=-\epsilon_{1j_{1}}$ we set $e(\ell_{1},j_{1},\epsilon_{1})=\emptyset$ again.
  • If we have $\epsilon_{1}=-$ and $\epsilon_{11}=-\epsilon_{1j_{1}}$ ,i.e. when we consider a term in the second line of Schwinger-Dyson equation (18), then we set:
    $e(1,j_{1},\epsilon_{1})$ $=((1,j)_{j\leq{j_{1}}},(1,j)_{j_{1}+1\leq j\leq m_{1}},(2,j)_{j\in[m_{2}]},...,(k,j)_{j\in[m_{k}]}),$
    $e_{rs}(1,j_{1},\epsilon_{1})$ $=((1,j)_{j\leq{j_{1}}},(2,j)_{j\leq m_{1}-j_{1}},(3,j)_{j\in[m_{2}]},...,(k+1,j)_{j\in[m_{k}]}))$
    $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})$ $=(((s_{1j},\epsilon_{1j})_{j\leq j_{1}}),(s_{1j},\epsilon_{1j})_{j_{1}+1\leq j\leq m_{1}},(s_{2j},\epsilon_{2j})_{j\leq m_{2}},...,(s_{kj},\epsilon_{kj})_{j\leq m_{k}})$
    We then construct $f_{1}$, $f_{1}^{rs}$, $e_{min}(\ell_{1},j_{1},\epsilon_{1})$ and $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})$ as we did in the previous case. Again for $\epsilon_{1}=-$ and $\epsilon_{11}=\epsilon_{1j_{1}}$ we set $e(\ell_{1},j_{1},\epsilon_{1})=\emptyset$.
We now consider $\ell_{1}\geq 2$ and $s_{11}=s_{\ell j_{1}}$. If $\epsilon_{1}=+1$ and $\epsilon_{11}=\epsilon_{1}\epsilon_{\ell j_{1}}=\epsilon_{\ell j_{1}}$ we refer to (19) and we have:
$\epsilon^{(1)}$ $:=-1$
$e(\ell_{1},j_{1},\epsilon_{1})$ $:=(((1,1),...,(1,m_{1}),(\ell_{1},j_{1}),(\ell_{1},j_{1}+1),...,(\ell_{1},m_{\ell_{1}}),(\ell_{1},1),...,(\ell_{1},j_{1}-1)),(2,j)_{j\in m_{2}},$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ...,(\ell_{1}-2,j)_{j\in[m_{\ell-2}]},(\ell_{1}-1,j)_{j\in[m_{\ell-1}]},(\ell_{1}+1,j)_{j\in[m_{\ell+1}]},...,(k,j)_{j\in[m_{k}]}),$
If $\epsilon_{1}=+$ and $\epsilon_{11}=-\epsilon_{\ell j_{1}}$ we set $e(\ell_{1},j_{1},\epsilon_{1})=\emptyset$. If $\epsilon_{1}=-1$ and $\epsilon_{11}=-\epsilon_{\ell_{1}j_{1}}$ we refer to (20) and we have:
$\epsilon^{(1)}$ $=+1$
$e(\ell_{1},j_{1},\epsilon_{1})$ $:=(((1,1),...,(1,m_{1}),(\ell_{1},j_{1}+1),(\ell_{1},j_{1}+2),...,(\ell_{1},m_{\ell_{1}}),(\ell_{1},1),...,(\ell_{1},j_{1})),(2,j)_{j\in m_{2}},$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ...,(\ell_{1}-2,j)_{j\in[m_{\ell-2}]},(\ell_{1}-1,j)_{j\in[m_{\ell-1}]},(\ell_{1}+1,j)_{j\in[m_{\ell+1}]},...,(k,j)_{j\in[m_{k}]}).$
Finally if $\epsilon_{1}=-$ and $\epsilon_{11}=\epsilon_{\ell_{1}j_{1}}$ we set $e(\ell_{1},j_{1},\epsilon_{1})=\emptyset$. It remains to construct the corresponding $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})$, $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})$ $f_{1}$, $f_{1}^{rs}$, $e_{min}(\ell_{1},j_{1},\epsilon_{1})$ and $\bar{e}(\ell_{1},j_{1},\epsilon_{1})$ as we did in the first case. In all the possible cases, if $e(\ell_{1},j_{1},\epsilon_{1})\neq\emptyset$, then there exists $k-1\leq k^{\prime}\leq k+1$, $m_{1}^{\prime},...,m_{k^{\prime}}^{\prime}$ and $(s_{\ell j}^{\prime},\epsilon_{\ell j}^{\prime})_{j\leq m_{\ell}^{\prime}}\in([d]\times\{+,-\})^{m_{\ell}}$ such that:
$\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})=((s_{1j}^{\prime},\epsilon_{1j}^{\prime})_{j\leq m_{1}^{\prime}},...,(s_{kj}^{\prime},\epsilon_{kj}^{\prime})_{j\leq m_{k^{\prime}}}^{\prime}).$
The data of $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\in[k]}$ minimal $k$-word and $(\ell_{1},j_{1},\epsilon_{1})$ pattern of length $1$ are the only data needed to recover $p_{1}$, $\epsilon^{(1)}$, $e(\ell_{1},j_{1},\epsilon_{1})$, $e_{rs}(\ell_{1},j_{1},\epsilon_{1})$ $e_{min}(\ell_{1},j_{1},\epsilon_{1})$, $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})$, $f_{1}$ and $f_{1}^{rs}$. We just described the first step of the algorithm applied to a minimal word $\mathcal{S}$ for a given pattern $(\ell_{1},j_{1},\epsilon_{1})$.
Definition 3.2.
Recall that $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\in[k]}$ is fixed and minimal and that we consider $m:=\sum_{\ell}m_{\ell}$. We fix $1\leq\ell_{0}\leq k$, $i\in[m_{\ell}]$ and a pattern $(\ell_{1},j_{1},\epsilon_{1})\in[k]\times[m]\times\{+,-\}$.
  • We say that the matrix in position $(\ell_{0},i)$ moves to position $(\ell,j)$ in the path $(\ell_{1},j_{1},\epsilon_{1})$ if $f_{1}(\ell_{0},i)=(\ell,j)$. We denote:
[…]
  • We say that the path $e(\ell_{1},j_{1},\epsilon_{1})$ terminates after one iteration if either $e(\ell_{1},j_{1},\epsilon_{1})=\emptyset$ or if denoting:
    $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})=((s_{1j}^{\prime},\epsilon_{1j}^{\prime})_{j\leq m_{1}^{\prime}},...,(s_{kj}^{\prime},\epsilon_{kj}^{\prime})_{j\leq m_{k}^{\prime}})$
    the word of first generation we have that for all $1\leq\ell\leq k$ the word $\mathcal{S}(\ell_{1},j_{1},\epsilon_{1})(\ell)=(s_{\ell j}^{\prime},\epsilon_{\ell j}^{\prime})_{j\leq m_{\ell}^{\prime}}$ is of trivial trace. It is equivalent to saying that the minimal word of first generation $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})$ is the empty word.
    We denote by $\mathcal{F}_{1}(\mathcal{S})\subset\{e(\ell_{1},j_{1},\epsilon_{1});\leavevmode\nobreak\ (\ell_{1},j_{1},\epsilon_{1})\in[k]\times[m]\times\{+,-\}\}$ the set of paths that terminates after one iteration.
Example 3.2.
We consider $\mathcal{S}=(\mathcal{S}(1),\mathcal{S}(2))$ where $\mathcal{S}(1)=((1,+),(1,+),(2,-),(3,-))$ and $\mathcal{S}(1)=((3,+),(2,+),(1,-),(1,-))$. We consider all the possible patterns of first generation $(\ell_{1},j_{1},\epsilon_{1})\in\{1,2\}\times[4]\times\{+,-\}$. In this example the patterns $(1,2,+)$, $(2,3,-)$ and $(2,4,-)$ are the patterns that give rise to paths $e(\cdot)$ different from the empty path. In other word the paths $e(1,1,\cdot)$, $e(1,2,-)$, $e(1,3,\cdot)$, $e(1,4,\cdot)$, $e(2,1,\cdot)$, $e(2,2,\cdot)$, $e(2,3,+)$ and $e(2,4,+)$ terminate after one iteration.
[…]
Figure 4: Result of the algorithm after one iteration of Schwinger-Dyson Equation
With the algorithm we are describing, we obtain a rooted tree structure that depends only on the original minimal word $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\in[k]}$. In Figure 4, the rooted tree in question is the tree obtained by removing the dashed edges (see Figure 5). In this example, the root $e_{0}$ is of degree $3$. In any case, for a given $m=\sum_{\ell}m_{\ell}$, and under the condition of having something other than an empty word, all nodes will have a degree bounded by $m$.
$e_{0}$$e(p_{3})$$e(p_{14})$$e(p_{16})$$p_{3}=(1,2,+)$$p_{14}=(2,3,-)$$p_{16}=(2,4,-)$
Figure 5: Tree structure for the first generation
Proposition 3.3.
Let $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\in[k]}$ a fixed original $k$-word. Following the previous construction we have that for all choices of $(\ell_{1},j_{1},\epsilon_{1})$:
$\sum_{\ell=1}^{k}m_{\ell}^{\prime}\leq m,$
where for all $1\leq\ell\leq k$, $m^{\prime}_{\ell}$ is the length of the $\ell$-th word in $\mathcal{S}_{min}(\ell_{1},j_{1},\epsilon_{1})$. Also in regards of Schwinger-Dyson equation (16) we have:
$\mathbb{E}^{\prime}(\mathcal{S})=\sum_{\ell_{1}=1}^{k}\sum_{j_{1}=1}^{m_{\ell}}\sum_{\epsilon_{1}\in\{+,-\}}\bar{e}(\ell_{1},j_{1},\epsilon_{1}).$
Example 3.3.
We continue with Example 3.2. Applying Schwinger-Dyson equation to
$\mathbb{E}^{\prime}(\mathcal{S}):=\mathbb{E}(\operatorname{Tr}[U(1)^{2}U(2)^{\ast}U(3)^{\ast}]\operatorname{Tr}[U(3)U(2){U(1)^{\ast}}^{2}])$
we obtain indeed:
$\mathbb{E}^{\prime}(\mathcal{S})=$ $-\frac{1}{N}\mathbb{E}(\operatorname{Tr}[U(1)]\operatorname{Tr}[U(1)U(2)^{\ast}U(3)^{\ast}]\operatorname{Tr}[U(3)U(2){U(1)^{\ast}}^{2}])$ (23)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{N}\mathbb{E}(\operatorname{Tr}[U(1)U(2)^{\ast}U(3)^{\ast}U(1)^{\ast}U(3)U(2)])$ (24)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{N}\mathbb{E}(\operatorname{Tr}[U(1)U(2)^{\ast}U(3)^{\ast}U(3)U(2)U(1)^{\ast}]).$ (25)
In fact, following the previous algorithm, we obtain that the term (23) corresponds to the term $\bar{e}(1,2,+)$, the term (24) to $\bar{e}(2,3,-)$ and finally the term (25) corresponds to $\bar{e}(2,4,-)$.
We now consider $n\geq 1$ and a pattern $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n+1]}\in(\operatorname{\mathbb{N}}\times\operatorname{\mathbb{N}}\times\{+,-\})^{n+1}$ of length $n+1$. We suppose that the data of $\mathcal{S}=((s_{\ell j},\epsilon_{\ell j})_{j\in[m_{\ell}]})_{\ell\in[k]}$ and $(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$ gave us the construction of $e(\cdot)$, $\mathcal{S}(\cdot)$, $e_{rs}(\cdot)$, $f_{b}$, $\mathcal{S}_{min}(\cdot)$, $e_{min}(\cdot)$, $f^{rs}_{b}$, $\epsilon^{(b)}$, $p_{b}$ and $\bar{e}(\cdot)$ of $(\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]}$ for all $1\leq b\leq n$. To ease notation in this second part of the algorithm, the pattern being fixed, for all $11\leq b\leq n+$ 1, we set $e_{b}=e((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]})$. Finally we suppose that for all $1\leq b\leq n$, denoting:
$\mathcal{S}_{min}((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]})=((s_{1j}^{(b)},\epsilon_{1j}^{(b)})_{j\in[m^{(b)}_{1}]},...,(s_{k_{b}j}^{(b)},\epsilon_{k_{b}j}^{(b)})_{j\in[m^{(b)}_{k_{b}}]})$
the minimal writing of the word of generation $b$, we have that:
$\sum_{\ell=1}^{k_{b}}m_{\ell}^{(b)}\leq m$
In particular if that is true for all patterns of length $1\leq b\leq n$ then one only has to consider $(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}\in([m]\times[m]\times\{+,-\})^{n}$. Under these assumptions we construct $e_{n+1}=e(\mathrm{P})$. If for some $1\leq b\leq n$, $e_{b}$ terminates after $b$ iterations we set $e_{n+1}=\emptyset$. Otherwise, we proceed as with the first generation. For instance we now denote:
$\mathcal{S}_{min}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})=((s_{1j}^{(n)},\epsilon_{1j}^{(n)})_{j\leq m_{1}^{(n)}},...,(s_{kj}^{(n)},\epsilon_{kj}^{(n)})_{j\leq m_{k_{n}}^{(n)}})$ (26)
the minimal word of the $n$-th generation. We also consider $\epsilon^{(n)}\in\{-1,+1\}$ and $p_{n}\in\mathbb{N}$ such that:
$\bar{e}_{n}=\epsilon^{(n)}\frac{N^{p_{n}}}{N^{n}}\mathbb{E}^{\prime}(\mathcal{S}_{min}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})).$
We recall that $\epsilon^{(n)}$ and $p_{n}$ depend in the pattern $(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$ and $\mathcal{S}$ the original word only. Once again if $\ell_{n+1}=1$ and $j_{n+1}=1$, or $\ell_{n+1}\geq k_{n}+1$, or $j_{n+1}\geq m_{\ell_{n+1}}+1$, or $s_{11}^{(n)}\neq s_{\ell_{n+1}j_{n+1}}^{(n)}$ we set $e_{n+1}=\emptyset$ and $\bar{e}_{n+1}=0$. Now we for instance suppose that $\ell_{n+1}\geq 2$, $\epsilon_{n+1}=+$ and $\epsilon_{11}^{(n)}=\epsilon_{n+1}\epsilon_{\ell_{n+1}j_{n+1}}^{(n)}=\epsilon_{\ell_{n+1}j_{n+1}}^{(n)}$, i.e when one considers a term of the third line of the Schwinger-Dyson equation (19). In this case we have:
$\epsilon^{(n+1)}$ $=-\epsilon^{(n)}$
$\text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e_{n+1}$ $:=(((1,1),...,(1,m_{1}^{(n)}),(\ell_{n+1},j_{n+1}),(\ell_{n+1},j_{n+1}+1),...,(\ell_{n+1},m_{\ell_{n+1}}^{(n)}),(\ell_{n+1},1),...,(\ell_{n+1},j_{n+1}-1))$
$(2,j)_{j\leq m_{2}^{(n)}},...,(\ell_{n+1}-1,j)_{j\leq m_{\ell_{n+1}-1}^{(n)}},(\ell_{n+1}+1,j)_{j\leq m_{\ell_{n+1}+1}^{(n)}},...,(k_{n},j)_{j\leq m_{k_{n}}^{(n)}})$
where all the notations refer to the notation chosen to explicit the minimal word of generation $n$ given by (26). We also construct the word of the $(n+1)$-th generation $\mathcal{S}_{n+1}=\mathcal{S}(\mathrm{P})$ as the $(k_{n}-1)$-word that verifies:
$\mathcal{S}_{n+1}(1)=$ $((s_{11}^{(n)},\epsilon_{11}^{(n)}),...,(s_{1m_{1}^{(n)}}^{(n)},\epsilon_{1m_{1}^{(n)}}^{(n)}),(s_{\ell_{n+1}j_{n+1}}^{(n)},\epsilon_{n+1}),(s_{\ell_{n+1}j_{n+1}},\epsilon_{n+1}),(s_{\ell_{n+1}j_{n+1}+1},\epsilon_{\ell_{n+1}j_{n+1}+1}),...,$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ...,(s_{\ell_{n+1}j_{n+1}-1},\epsilon_{\ell_{n+1}j_{n+1}-1})),$
$\mathcal{S}_{n+1}(\ell)=$ $(s_{\ell j},\epsilon_{\ell j})_{j\leq m_{\ell}^{(n)}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ 2\leq\ell\leq\ell_{n+1}-1$
$\mathcal{S}_{n+1}(\ell-1)=$ $(s_{\ell j},\epsilon_{\ell j})_{j\leq m_{\ell}^{(n)}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{n+1}+1\leq\ell\leq k_{n}.$
Finally, we have the $n+1$-th tracking function:
$f_{n+1}:$ $(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$ $\rightarrow(\mathbb{N}\times\mathbb{N})^{(\mathbb{N}\times\mathbb{N})}$
$(\ell,j)$ $\mapsto(\ell,j)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ \leavevmode\nobreak\ 1\leq\ell\leq\ell_{n+1}-1.$
$(\ell_{n+1},j)$ $\mapsto(1,m_{1}+1+j-j_{1})$
$(\ell,j)$ $\mapsto(\ell-1,j)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{n+1}+1\leq\ell\leq k_{n}.$
We now may recover the corresponding word $\mathcal{S}((\ell_{p},j_{p},\epsilon_{p})_{p\in[n+1]})$, its corresponding minimal writing, the rescaled path $e_{rs}$, minimal path $e_{min}$ and rescaled tracking function $f_{n+1}^{rs}$. Finally we consider $t_{n+1}$ the number of $11\leq\ell\leq k_{n}-$ 1 such that $\mathcal{S}_{n+1}(\ell)$ is of trivial trace and set $p_{n+1}=p_{n}+t_{n+1}$. We now have recovered the mean of paths $(\ell_{p},j_{p},\epsilon_{p})_{p\in[n+1]}$: We proceed similarly for the other cases of $(\ell_{n+1},j_{n+1},\epsilon_{n+1})$. We may now introduce the following Definition of terminating terms after $n$ iterations of Schwinger-Dyson equations.
$\bar{e}((\ell_{p},j_{p},\epsilon_{p})_{p\in[n+1]})=\epsilon^{(n+1)}\frac{N^{p_{n+1}}}{N^{n+1}}\mathbb{E}^{\prime}(\mathcal{S}_{min}((\ell_{p},j_{p},\epsilon_{p})_{p\in[n+1]})).$
We proceed similarly for the other cases of $(\ell_{n+1},j_{n+1},\epsilon_{n+1})$. We may now introduce the following Definition of terminating terms after $n$ iterations of Schwinger-Dyson equations.
Definition 3.4.
We say that the path $e((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})$ terminates after $n$ iterations if either $e((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})=\emptyset$ or if denoting:
$\mathcal{S}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})=((s_{1j}^{(n)},\epsilon_{1j}^{(n)})_{j\in[m_{1}^{(n)}]},...,(s_{kj}^{(n)},\epsilon_{k_{n}j}^{(n)})_{j\in[m_{k_{n}}^{(n)}]})$
the word of generation $n$, we have that for all $1\leq\ell\leq k_{n}$ we have that the $1$-word $\mathcal{S}(\ell)=(s_{\ell j}^{(n)},\epsilon_{\ell j}^{(n)})_{j\leq m_{\ell}^{(n)}}$ is of trivial trace. We denote by $\mathcal{F}_{n}(\mathcal{S})\subset\{e((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})\in(\mathbb{N}\times\mathbb{N}\times\{+,-\})^{n}\}$ the set of paths that terminates after $n$ iteration.
Proposition 3.5.
We consider $\mathcal{S}$ any original fixed minimal $k$-word. Let $n\geq 1$ be an integer and $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}\in([m]\times[m]\times\{+,-\})^{n}$ a pattern. Using the notations of the previous algorithm applied to $\mathcal{S}$ and $\mathrm{P}$ we have:
$\sum_{\ell=1}^{k_{n}}m_{\ell}^{(n)}\leq m$
where we used the notations given for the minimal writing of the $n$-th generation word obtained by the algorithm (26). Also we have:
$\mathbb{E}^{\prime}(\mathcal{S})=\sum_{b=1}^{n-1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{e\in\mathcal{F}_{b}(\mathcal{S})}\bar{e}+\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}\in([m]\times[m]\times\{+,-\})^{n}}\bar{e}(\mathrm{P}).$ (28)
Finally the number of paths $e(\mathrm{P})$ with $\mathrm{P}\in([m]\times[m]\times\{+,-\})^{n}$ that did not terminates before the $n$-th iteration is bounded by $(m-1)^{n}$.
Proof.
For all $n\geq 1$ and all patterns $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b^{\in}[n]}$, the construction of $\mathcal{S}(\mathrm{P})$ is such that the total number of matrices appearing in $\mathcal{S}(\mathrm{P})$ is exactly given by the total number of matrices in $\mathcal{S}_{min}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n-1]})$. Thus the total number of matrices after simplification, i.e. after replacing the subwords by their minimal writing, is bounded by the original number of matrices. The second statement is given by induction on $n\geq 1$. In fact, the initiation is given by Proposition 3.3. If we now assume that the formula (28) is verified for some $n\geq 1$, one has to apply the Schwinger-Dyson equation (16) to each term $\bar{e}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})$ to obtain the same formula for $n+1$. Recall that we denote by
$\mathcal{S}_{min}((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]})=((s_{1j}^{(b)},\epsilon_{1j}^{(b)})_{j\in[m^{(b)}_{1}]},...,(s_{k_{b}j}^{(b)},\epsilon_{k_{b}j}^{(b)})_{j\in[m^{(b)}_{k_{b}}]})$
the corresponding minimal word of generation $b$. To have a non-empty path of $(b+1)$-th generation one has to consider $(\ell_{b+1},j_{b+1},\epsilon_{b+1})$ such that $(\ell_{b+1},j_{b+1})\neq(1,1)$ and $s_{11}^{(b)}=s_{\ell_{b+1},j_{b+1}}^{(b)}$. Also if we have $\epsilon_{b+1}=+$, then the corresponding path $e_{b+1}$ of generation $b+1$ is non-empty if and only if $\epsilon_{\ell_{b+1}j_{b+1}}^{(b)}=\epsilon_{11}^{(b)}$. If we have $\epsilon_{b+1}=-$, the corresponding path of generation $b+1$ is non-empty if and only if $\epsilon_{\ell_{b+1}j_{b+1}}^{(b)}=\epsilon_{11}^{(b)}$. Therefore for all $(\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b+1]}$, if $e(\ell_{1},j_{1},\epsilon_{1},...,\ell_{b+1},j_{b+1},\epsilon_{b+1})\neq\emptyset$ then $e(\ell_{1},j_{1},\epsilon_{1},...,\ell_{b+1},j_{b+1},-\epsilon_{b+1})=\emptyset$. Therefore for $(\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]}$ fixed we have:
|$|\{(\ell_{b+1},j_{b+1},\epsilon_{b+1})\in[m]\times[m]\times\{+,-\};\leavevmode\nobreak\ e((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b+1]})\neq\emptyset\}|$ $\leq\sum_{\ell,j}\mathds{1}(s_{11}^{(b)}=s_{\ell j}^{(b)},\leavevmode\nobreak\ (1,1)\neq(\ell,j))$
$\leq m-1.$
From this last inequality we obtain the proposition’s last assertion by induction on $1\leq b\leq n$. ∎
3.2 Writing the iterations of Schwinger-Dyson and convergence of series
With the previous proposition, and for some values of $m\geq 1$, we can now express the expectation of the traces as a series of terminated terms given by the algorithm.
Proposition 3.6.
Let:
$\mathcal{S}=((s_{1j},\epsilon_{1j})_{j\in[m_{1}]},...,(s_{kj},\epsilon_{kj})_{j\in[m_{k}]})$
be any original minimal $k$-word such that $m:=\sum_{\ell}m_{\ell}<N^{2/3}+1$. We apply the previous algorithm and denote by $\mathcal{F}_{n}(\mathcal{S})$ the set of finishing path after $n$-iteration. We have that:
$\sum_{n}\sum_{e\in\mathcal{F}_{n}(\mathcal{S})}\bar{e}$
is convergent and furthermore considering the expectation given by (15) we have for $N$ large enough:
$\mathbb{E}^{\prime}(\mathcal{S})=\sum_{n=0}^{\infty}\sum_{e\in\mathcal{F}_{n}(\mathcal{S})}\bar{e}.$
The main argument to prove the previous proposition is the following lemma.
Lemma 3.7.
For any original minimal $k$-word denoted:
$\mathcal{S}=((s_{1j},\epsilon_{1j})_{j\in[m_{1}]},...,(s_{kj},\epsilon_{kj})_{j\in[m_{k}]})$
using the same notations than in Proposition 3.6 we have for all $n\geq 1$ integer:
$|\mathcal{F}_{n}(\mathcal{S})|$ $\leq(m-1)^{n}$
$\forall e\in\mathcal{F}_{n}(\mathcal{S})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |\bar{e}|$ $\leq N^{k-2/3n}.$
The proof of this lemma follows the proof given by Hastings [1, Part C.].
Proof.
Let $e=e((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})\in\mathcal{F}_{n}(\mathcal{S})$ such that $\bar{e}\neq 0$. Recall that for all $1\leq b\leq n$ we denote by $k_{b}$ the integer such that the minimal writing $\mathcal{S}_{min}((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{{b^{\prime}}\in[b]})$ of generation $b$ is a $k_{b}$ word, $p_{b}\in\mathbb{N}$ such that:
$|\bar{e}((\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]})|=\frac{N^{p_{b}}}{N^{b}}|\mathbb{E}^{\prime}(\mathcal{S}_{min}((\ell_{b}^{\prime},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]}))|$
where $\mathbb{E}^{\prime}$ defined by Equation (15). Although we set $p=p_{n}$. We denote by $q:=|\{b\leq n;\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{b}=1\}|$ the number of times we have considered a path that gives a term in the first (17) or second (18) line of the Schwinger-Dyson equations. For all $1\leq b\leq n$ such that $\ell_{b}=1$ we necessarily have $k_{b}-k_{b-1}=1$. Indeed, since we assume at each step that we are applying our algorithm to minimal words that are not empty, cutting a word in half cannot produce a word of trivial trace and therefore the minimal word of the $b$-th generation is a $(k_{b-1}+1)$-word. Also the integers $1\leq b\leq n$ such that $k_{b}-k_{b-1}=-2$ are exactly the terms where $p_{b}=p_{b-1}+1$. Finally, $e$ being a finishing term, we have $k_{n}=0$, $k_{0}=k$ and therefore:
$k_{n}=0$ $=k_{0}+\sum_{b=1}^{n}k_{b}-k_{b-1}$
$=k+\sum_{b=1}^{n}\mathds{1}(\ell_{b}=1)-\sum_{b=1}^{n}(\mathds{1}(\ell_{b}\neq 1)+\mathds{1}(p_{b}=p_{b-1}+1))$
$=k+q-(n-q)-p.$
The minimality of the words to which we apply the Schwinger-Dyson equation also implies that the words of trivial traces come from the iterations $\{1\leq b\leq n;\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{b}\neq 1\}$, i.e. when we consider terms from the third (19) or fourth (20) line. In particular we have:
$p\leq n-q$
and therefore:
$p\leq\frac{k+n}{3}.$
Finally, the argument for the bound on $|\mathcal{F}_{n}(\mathcal{S})|$ is exactly the one given to prove the last assertion of Proposition 3.5, noting that $\mathcal{F}_{n}$ is included in the number of non-empty terms. ∎
We now give the proof of the previous proposition.
Proof of Proposition 3.6.
The convergence of the series is given by the fact that we have:
$\left|\sum_{e\in\mathcal{F}_{n}(\mathcal{S})}\bar{e}\right|\leq N^{k}\left(\frac{m-1}{N^{2/3}}\right)^{n}$
where we have $\frac{m-1}{N^{2/3}}<1$. For $n\geq 1$ we denote:
$r_{n}:=\sum_{\mathrm{P}\in([m]\times[m]\times\{+,-\})^{n}}\bar{e}(\mathrm{P}).$
We notice that for all $n\geq 1$ and for all $\mathrm{P}$, using the notation of the algorithm we have that $\sum_{\ell=1}^{k_{n}}m^{(n)}_{\ell}\leq m$ and for all $1\leq\ell\leq k_{n}$ we have:
$|\operatorname{Tr}[U(s_{\ell 1}^{(n)})^{\epsilon_{\ell 1}^{(n)}}\cdots U(s_{\ell m^{(n)}_{\ell}}^{(n)})^{\epsilon_{\ell m^{(n)}_{\ell}}^{(n)}}]|\leq N$
and therefore:
$|r_{n}|$ $\leq\sum_{\mathrm{P}\in([m]\times[m]\times\{+,-\})^{n}}|\bar{e}(\mathrm{P})|$
$\leq(m-1)^{n}\frac{N^{m}}{N^{n}}.$
For $N$ large enough we have $(m-1)/N<1$ and therefore $|r_{n}|\underset{n\rightarrow\infty}{\longrightarrow}0$. To conclude we apply Proposition 3.5 and we have:
$|r_{n}|=\left|\mathbb{E}_{0}(\mathcal{S})-\sum_{b=1}^{n-1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{e\in\mathcal{F}_{b}(\mathcal{S})}\bar{e}\right|.$
4 Computations of traces
Throughout this section we consider $\mathcal{S}=((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))$ a minimal $1$-word of length $m$ and we want to compute:
$\mathbb{E}^{\prime}(\tilde{\mathcal{S}})=\mathbb{E}_{0}(\mathcal{S})=\mathbb{E}(\operatorname{Tr}[\mathcal{U}(\mathcal{S})]\operatorname{Tr}[\mathcal{U}(\mathcal{S})^{\ast}])$
where we set $\tilde{\mathcal{S}}=(\tilde{\mathcal{S}}(1),\tilde{\mathcal{S}}(2))$ with $\tilde{\mathcal{S}}(1)=\mathcal{S}$ and $\tilde{\mathcal{S}}(2)=((s_{m},-\epsilon_{m}),...,(s_{1},-\epsilon_{1}))$ (see Equations (15) and (4) for definitions of $\mathbb{E}^{\prime}(\cdot)$ and $\mathbb{E}_{0}(\cdot)$). Since the previous definition of $\tilde{\mathcal{S}}$ depends only $\mathcal{S}$ we abuse notations and use the notation $\mathcal{S}$ to refer to the $2$-word $\tilde{\mathcal{S}}$.
Definition 4.1.
We fix $\ell_{0}\in\{1,2\}$, $i\in[m]$ and a pattern $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}\in([m]\times[m]\times\{+,-\})^{n}$. We say that the matrix in position $(\ell_{0},i)$ moves to position $(\ell,j)$ in the path $e(\mathrm{P})$ at the $n$-th iteration if $f_{n}\circ f_{n-1}^{rs}\circ\cdots\circ f_{1}^{rs}(\ell_{0},i)=(\ell,j)$. We denote:
[…]
where for all $11\leq b\leq n-$ 1 we set $(\ell_{b}(i),j_{b}(i))=f^{rs}_{b}(\ell_{b-1}(i),j_{b-1}(i))$ with convention $(\ell_{0}(i),j_{0}(i)):=(1,m+1-i)$. Similarly we denote:
[…]
where for all $11\leq b\leq n-$ 1, $(\ell_{b}^{\prime}(i),j_{b}^{\prime}(i))=f^{rs}_{b}(\ell_{b-1}^{\prime}(i),j_{b-1}^{\prime}(i))$ with convention $(\ell_{0}^{\prime}(i),j_{0}^{\prime}(i)):=(2,i)$.
4.1 Rung cancellations of matrices
In this section we analyze the terminating non-zero terms of generation $n$, verifying that the matrix originally at position $(\ell_{0},i)$ was never used before the cancellation, and is then canceled against its original inverse in the second trace (i.e., the matrix originally at position (¯$(\bar{\ell}_{0},m+1-i)$ with ¯$\bar{\ell}_{0}\neq\ell_{0}$).
Definition 4.2.
[…]
$f_{n}^{rs}(\ell_{n-1}(i),j_{n-1}(i))\notin\{(\ell_{n+1},j_{n+1}),(1,1)\}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{n}^{rs}(\ell_{n-1}^{\prime}(i),j_{n-1}^{\prime}(i))\notin\{(\ell_{n+1},j_{n+1}),(1,1)\}.$
If we suppose $e\in\mathcal{F}_{n+1}(\mathcal{S})$, we say that we have a rung cancellation of matrix $i$ if there exists $n_{i}\leq n+1$ such that for all $b\leq n_{i}$, the matrix $i$ is trivially moved under the $b$-th iteration and if $f_{n_{i}}(\ell_{n_{i}-1}^{\prime}(i),j_{n_{i}-1}^{\prime}(i))=f_{n_{i}}(\ell_{n_{i}-1}(i),j_{n_{i}-1}(i))\pm 1$ (where we take it as a convention that $(\ell^{\prime},j)\pm 1=(\ell^{\prime},j^{\prime}\pm 1)$ and $(\ell,m_{\ell}+1)=(\ell,1)$ and $(\ell,1-1)=(\ell,m_{\ell})$).
Remark 4.1.
One can interpret the previous definition as follows: the matrix $i$ is trivially moved under the $n$-th iteration if the term we consider after applying the Schwinger-Dyson equation does not originate from a term of the form $TU(s_{\ell i})$ (see [1] and proof of Proposition 2.1). We have a rung cancellation of the matrix $i$ if in all the previous paths we never crossed the matrices originally at positions $(1,2m+1-i)$ and $(2,i)$ and at some point they canceled each other.
We will now describe the terms finishing at the $n$-th iteration and having a rung cancellation of matrix $i$ for any $i\in[m]$. Let $X=U(d+1)$ Haar distributed and independent from $(U(s))_{s\in[d]}$. For $1\leq i\leq m$ we define:
$\mathbb{E}_{0}^{i}(\mathcal{S})$ $:=\mathbb{E}(\operatorname{Tr}[U(s_{m})^{\epsilon_{m}}\cdots U(s_{i+1})^{\epsilon_{i+1}}XU(s_{i-1})^{\epsilon_{i-1}}\cdots U(s_{1})^{\epsilon_{1}}]\times$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{Tr}[U(s_{1})^{-\epsilon_{1}}\cdots U(s_{i-1})^{-\epsilon_{i-1}}X^{\ast}U(s_{i+1})^{-\epsilon_{i+1}}\cdots U(s_{2m})^{\ast}]).$
We then set:
$e_{0}^{i}=e_{0}$
with $e_{0}$ as defined in the beginning of Section 3.1. We apply the previous algorithm to $e^{i}_{0}$ considering as fixed original word:
$\mathcal{S}^{i}:=((s_{1},\epsilon_{1}),...,(s_{i-1},\epsilon_{i-1}),(d+1,+),(s_{i+1},\epsilon_{i+1}),...,(s_{m},\epsilon_{m})).$
In particular we have $\mathcal{S}^{i}$ also minimal.
Proposition 4.3.
Let $\mathrm{P}=(\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]}\in(\mathbb{N}\times\mathbb{N}\times\{+,-\})^{n}$ be a pattern. If the path $e:=e((\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]})$ terminates at the $n$-th iteration and has a rung cancellation of matrix $i$ then $e^{i}:=e^{i}((\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]})=e((\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]})$ and their mean are equal. Conversely if $e^{i}((\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]})$ is a term that terminates at the $n$-iteration when applying the algorithm to $\mathcal{S}^{i}$, then $e$ terminates at the $n$-th iteration, has a rung cancellation of matrix $i$ and has same mean than $e^{i}$. In particular we have for all $n\in\mathbb{N}$:
$\mathcal{F}_{n}(\mathcal{S}^{i})\subset\mathcal{F}_{n}(\mathcal{S}).$
We then set for all $n\geq 1$:
$\mathcal{R}_{n}(\mathcal{S}):=\mathcal{F}_{n}(\mathcal{S})\backslash\bigcup_{i\in[m]}\mathcal{F}_{n}(\mathcal{S}^{i}).$
Definition 4.4.
Let $1\leq i_{1}<\cdots<i_{q}\leq m$ be elements of $[m]$ and $e=e((\ell_{b},\ell_{b},\epsilon_{b})_{b\in[n]})\neq\emptyset$ be a path that terminates at the $n$-th iteration. We say that there is a rung cancellation of matrices $\vec{i}=(i_{1},...,i_{q})$ if for all $1\leq t\leq q$ there is a rung cancellation of matrix $i_{t}$.
We fix $1\leq i_{1}<\cdots<i_{q}\leq m$ and we generalize the previous description of the terms with a rung cancellation of matrices $\vec{i}=(i_{1},...,i_{q})$. We consider $(X_{1},...,X_{q})=(U(d+1),...,U(d+q))$ such that $(U(s))_{s\in[d+q]}$ are Haar distributed independent unitaries. We denote:
$\mathbb{E}_{0}^{\vec{i}}(\mathcal{S})=\mathbb{E}(\operatorname{Tr}[U(s_{m})^{\epsilon_{m}}$ $\cdots U(s_{i_{t}+1})^{\epsilon_{i_{t}+1}}X_{t}U(s_{i_{t}-1})^{\epsilon_{i_{t}-1}}\cdots U(s_{1})^{\epsilon_{1}}]\times$
$\operatorname{Tr}[U(s_{1})^{-\epsilon_{1}}\cdots U(s_{i_{t}-1})^{-\epsilon_{i_{t}-1}}X_{t}^{\ast}U(s_{i_{t}+1})^{-\epsilon_{i_{t}+1}}\cdots U(s_{m}^{-\epsilon_{m}})]).$
We set as for the case $q=1$, $e^{\vec{i}}_{0}=e_{0}$ and we then apply the previous algorithm to $e^{\vec{i}}_{0}$ by considering the corresponding fixed original word of the algorithm to be:
$\mathcal{S}^{\vec{i}}_{0}$ $=((s_{1},\epsilon_{1}),...,(s_{i_{1}-1},\epsilon_{i_{1}-1}),(d+1,+),(s_{i_{1}+1},\epsilon_{i_{1}+1}),...$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ...,(s_{i_{t}-1},\epsilon_{i_{t}-1}),(d+t,+),(s_{i_{t}+1},\epsilon_{i_{t}-1}),...,(s_{i_{q}-1},\epsilon_{i_{q}-1}),(d+q,+),(s_{i_{q}+1}\epsilon_{i_{q}+1}),...,(s_{m},\epsilon_{m})).$
For all $n\in\mathbb{N}$ we denote by $\mathcal{F}_{n}(\mathcal{S}^{\vec{i}})$ the set of terms $e^{\vec{i}}((\ell_{b},j_{b},\epsilon_{b})_{b\in[n]})$ that terminates at the $n$-th iteration.
Proposition 4.5.
Let $1\leq q\leq m$ and $1\leq i_{1}<\cdots<i_{q}\leq m$ be fixed. The Proposition 4.3 holds when we replace $i$ by $\vec{i}$. Besides we also have:
  • $\mathcal{F}_{n}(\mathcal{S}^{\vec{i}})=\bigcap_{t=1}^{q}\mathcal{F}_{n}(\mathcal{S}^{{i_{t}}})$
  • $\mathbb{E}_{0}^{\vec{i}}(\mathcal{S})=1$.
Proof.
The first point follows by induction, noticing that denoting $\vec{i}_{t}=(i_{1},...,i_{t})$ and applying the algorithm to $\mathcal{S}^{\vec{i}_{t}}$ then the terms with rung cancellation of matrices $(i_{1},...,i_{q})$ in the algorithm applied to $\mathcal{S}$ are the terms with rung cancellations of matrix $i_{q}$ in the algorithm applied to $\mathcal{S}^{\vec{i}_{q-1}}$.
For the second point we start with the case where $q=1$:
$\mathbb{E}_{0}^{i}(\mathcal{S})$ $=\mathbb{E}(\operatorname{Tr}[XU(s_{i-1})^{\epsilon_{i-1}}\cdots U(s_{1})^{\epsilon_{1}}U(s_{m})^{\epsilon_{m}}\cdots U(s_{i+1})^{\epsilon_{i+1}}]\times$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{Tr}[X^{\ast}U(s_{i+1})^{-\epsilon_{i+1}}\cdots U(s_{m})^{-\epsilon_{m}}U(s_{1})^{-\epsilon_{1}}\cdots U(s_{i-1})^{-\epsilon_{i-1}}])$
$=\frac{1}{N}\operatorname{Tr}[U(s_{i-1})^{\epsilon_{i-1}}\cdots U(s_{1})^{\epsilon_{1}}U(s_{m})^{\epsilon_{m}}\cdots U(s_{i+1})^{\epsilon_{i+1}}U(s_{i+1})^{-\epsilon_{i+1}}\cdots U(s_{m})^{-\epsilon_{m}}U(s_{1})^{-\epsilon_{1}}\cdots U(s_{i-1})^{-\epsilon_{i-1}}]$
$=1$
where we applied the Schwinger-Dyson equation between the first and second equality. The general case comes from induction, noticing that:
$\mathbb{E}_{0}^{\vec{i}}(\mathcal{S})=\mathbb{E}_{0}^{i_{q}}(\mathcal{S}^{\vec{i}_{q-1}})=\mathbb{E}_{0}^{i_{q}}((s_{1},\epsilon_{1}),...,(s_{i_{t}-1},\epsilon_{i_{t}-1}),(d+t,+),(s_{i_{t}+1},\epsilon_{i_{t}+1}),...,(s_{m},\epsilon_{m})).$
We may now give the following lemma and first computation of $\mathbb{E}_{0}(\mathcal{S})$.
Lemma 4.6.
Let $\mathcal{S}=((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))\in([d]\times\{+,-\})^{m}$ with $m<\frac{1}{2}(N^{2/3}+1)$. There exists a unique $\mathcal{S}^{\prime}=((s^{\prime}_{1},\epsilon_{1}^{\prime}),...,(s^{\prime}_{m^{\prime}},\epsilon_{m^{\prime}}))$ minimal writing of $((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))\in([d]\times\{+,-\})^{m}$ and we have:
$\mathbb{E}_{0}(\mathcal{S})=1+\sum_{n=0}^{\infty}\sum_{e\in\mathcal{R}_{n}(\mathcal{S}^{\prime})}\bar{e}.$
Also for all integer $n$ we have $\mathcal{F}_{n}(\mathcal{S})=\mathcal{F}_{n}(\mathcal{S}^{\prime})$ and $\mathcal{R}_{n}(\mathcal{S})=\mathcal{R}_{n}(\mathcal{S}^{\prime})$.
Proof.
We have:
$\mathbb{E}_{0}(\mathcal{S})=\mathbb{E}_{0}(\mathcal{S}^{\prime})$
therefore we consider directly $\mathcal{S}$ minimal word and we have:
$\mathbb{E}_{0}(\mathcal{S})=\sum_{n=0}^{\infty}\sum_{e\in\mathcal{R}_{n}(\mathcal{S})}\bar{e}+\underbrace{\sum_{n=0}^{\infty}\sum_{e\in\bigcup_{i=1}^{m}\mathcal{F}_{n}(\mathcal{S}^{i})}\bar{e}}_{\mathbb{E}_{0}^{r}}$
where we applied Proposition 3.6 for the first equality. We want to show that $\mathbb{E}_{0}^{r}=1$. Indeed:
$\mathbb{E}_{0}^{r}$ $=\sum_{n=0}^{\infty}\sum_{e\in\bigcup_{i=1}^{m}\mathcal{F}_{n}(\mathcal{S}^{i})}\bar{e}=\sum_{n=0}^{\infty}\sum_{q=1}^{m}(-1)^{q-1}\sum_{i_{1}<...<i_{q}}\sum_{e\in\mathcal{F}_{n}(\mathcal{S}^{\vec{i}})}\bar{e}$
$=\sum_{n=0}^{\infty}\sum_{e\in\mathcal{F}_{n}(\mathcal{S}^{\vec{i}})}\bar{e}=\sum_{q=1}^{m}(-1)^{q-1}\sum_{i_{1}<...<i_{q}}\mathbb{E}_{0}^{\vec{i}}(\mathcal{S}^{\prime})$
$=\sum_{q=1}^{m}(-1)^{q-1}\binom{m}{q}=1,$
where in the first line of calculation we applied the inclusion-exclusion principle to the union $\bigcup_{i\in[m]}\mathcal{F}_{n}(\mathcal{S}^{i})$ and the first point of Proposition 4.5. ∎
All that remains is to make conditions explicit on the original word $((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))\in([d]\times\{+,-\})^{m}$ to have terms without rung cancellation and then use these conditions to compute either $\mathbb{E}_{1}$ or $\mathbb{E}_{2}$ (defined by Equations (3) and (8) respectively ). To do this, we have to take into account the dependence for paths in the original word. From now on, we will specify the dependence in $\mathcal{S}$ for all sequences obtained by the algorithm in Section 3. For example, for $\mathcal{S}$ a minimal word and $\mathrm{P}$ a pattern of length $n$, we will denote by $e(\mathcal{S},\mathrm{P})$ the path of generation $n$ obtained by applying the algorithm to $\mathcal{S}$ and $\mathrm{P}$, where before we would have denoted $e(\mathrm{P})$ without confusion.
Definition 4.7.
For all $m,p\geq 1$ integers and we denote:
  • $\mathcal{W}(m)$ the set of minimal words $\mathcal{S}=((s_{1},\epsilon_{1}),...,(s_{m},\epsilon_{m}))$ of length $m$,
  • $\mathcal{W}^{\prime}(m)$ the subset of minimal words of length $m$ such that there exists a non-empty finishing path with no rung cancellation of matrix.
  • $\mathcal{W}_{1}(p)$ the set of minimal words $\mathcal{S}=((s_{1},+),(s_{2},-),...,(s_{2p-1},+),(s_{2p},-))$ of length $p$, and likewise $\mathcal{W}_{1}^{\prime}$ the subset of $\mathcal{W}_{1}(p)$ of minimal word that give rise to non-empty finishing path with no rung cancellation.
Let $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$ be a pattern of length $n$. We set $\mathcal{W}^{\prime}(m,\mathrm{P})\subset\mathcal{W}^{\prime}(m)$ (resp. $\mathcal{W}_{1}^{\prime}(p,\mathrm{P})\subset\mathcal{W}^{\prime}_{1}(p)$) the set of $\mathcal{S}\in\mathcal{W}^{\prime}(m)$ (resp. $\mathcal{S}\in\mathcal{W}^{\prime}_{1}(p)$) such that $e(\mathcal{S},\mathrm{P})\neq\emptyset$ terminates at the $n$-th generation. We denote by $[\mathcal{S}]_{\mathrm{P}}$ the set of $\mathcal{S}^{\prime}\in\mathcal{W}^{\prime}(m,\mathrm{P})$ (resp. $\mathcal{S}^{\prime}\in\mathcal{W}_{1}^{\prime}(p,\mathrm{P})$) of minimal words of same length and such that for all $1\leq b\leq n$ we have:
$f_{b}(\mathcal{S}^{\prime},\mathrm{P})(\cdot)=f_{b}(\mathcal{S},\mathrm{P})(\cdot)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{b}^{rs}(\mathcal{S}^{\prime},\mathrm{P})(\cdot)=f_{b}^{rs}(\mathcal{S},\mathrm{P})(\cdot).$
where $f_{b}(\mathcal{S},\mathrm{P})(\cdot),f_{b}^{rs}(\mathcal{S},\mathrm{P})(\cdot)$ are the tracking and rescaled tracking functions defined in (3.1). This last definition constitutes an equivalence relation that we denote by $\sim_{\mathrm{P}}$.
Lemma 4.8.
Let $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}\in([2m]\times[2m]\times\{+,-\})^{n}$ be a pattern and $\mathcal{S}\in\mathcal{W}^{\prime}(m,\mathrm{P})$. Then there exists $f:[m]\rightarrow[m]$ with no fixed point and such that for all $\mathcal{S}^{\prime}=((s_{1}^{\prime},\epsilon_{1}^{\prime}),...,(s_{m}^{\prime},\epsilon^{\prime}_{m}))\in[\mathcal{S}]_{\mathrm{P}}$ we have for all $1\leq i\leq m$:
$s_{f(i)}^{\prime}=s_{i}^{\prime}.$
Likewise for $\mathcal{S}\in\mathcal{W}^{\prime}_{1}(p,\mathrm{P})$, there exists $f:[2p]\rightarrow[2p]$ such that for all $\mathcal{S}^{\prime}=((s_{1}^{\prime},+),(s_{2}^{\prime},-),...,(s_{2p-1}^{\prime},+),(s_{2p},-))$ and for all $1\leq i\leq 2p$ we have:
$s_{f(i)}^{\prime}=s_{i}^{\prime}.$
Proof.
The proof is the same in both cases of the previous lemma, therefore we prove only the first case. For all $1\leq b\leq n$ we consider $e_{b}:=e(\mathcal{S},\mathrm{P}_{b})$ where $\mathrm{P}_{b}=(\ell_{b^{\prime}},j_{b^{\prime}},\epsilon_{b^{\prime}})_{b^{\prime}\in[b]}$. We construct the function $f:[m]\rightarrow[m]$ as follows. First we notice that it is sufficient to show that for all $i\in[m]$ there exists $j\neq i$ such that for all $\mathcal{S}^{\prime}=((s_{k}^{\prime},\epsilon_{k}^{\prime})_{k\in[m]})\sim_{\mathrm{P}}\mathcal{S}$ we have $s_{i}^{\prime}=s_{j}^{\prime}$. Let now $i\in[m]$ be any index. We first suppose that there exists $11\leq b\leq n-$ 1 such that $(\ell_{b}(i),j_{b}(i))\in\{(1,1),(\ell_{b},j_{b})\}$ or $(\ell_{b}^{\prime}(i),j_{b}^{\prime}(i))\in\{(1,1),(\ell_{b+1},j_{b+1})\}$. In other words we suppose that for some iteration, the matrix $i$ is not trivially moved in the path $e(\mathcal{S},\mathrm{P})$ (cf Definition 4.2). We consider the smallest $11\leq b\leq n-$ 1 verifying this condition. If $(\ell_{b}(i),j_{b}(i))=(1,1)$ and since $e_{b+1}\neq\emptyset$, then there exists $j\in[m]$ such that $(\ell_{b+1},j_{b+1})\in\{(\ell_{b}(j),j_{b}(j)),(\ell_{b}^{\prime}(j),j_{b}^{\prime}(j))\}$ and therefore $s_{i}=s_{j}$. In the case where $(\ell_{b+1},j_{b+1})=(\ell_{b}(j),j_{b}(j))$ it means is we have the paths:
[…]
By uniqueness of the paths we have $m+1-i\neq m+1-j$ and therefore $i\neq j$. If now we have $(\ell_{b+1},j_{b+1})=(\ell_{b}^{\prime}(j),j_{b}^{\prime}(j))$, we have the paths:
[…]
Since we assumed $b$ the first time that the matrix $i$ is not trivially moved and that we have no rung cancellation of the matrix $i$, then $i\neq j$. In the cases where $(\ell_{b}(i),j_{b}(i))=(\ell_{b+1},j_{b+1})$ or $(\ell_{b}^{\prime}(i),j_{b}^{\prime}(i))\in\{(1,1),(\ell_{b+1},j_{b+1})\}$ we similarly show the existence of $j\neq i$ such that $s_{i}=s_{j}$.
Now if we suppose the matrix $i$ is never trivially moved. Still the term $e_{n}$ being of trivial traces, there exists $n_{i}\leq n$ such that:
$f_{n_{i}}(\ell_{n_{i}-1}(i),j_{n_{i}-1}(i))$ $\neq(0,0)$
$\text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{n_{i}}^{rs}(\ell_{n_{i}-1}(i),j_{n_{i}-1}(i))$ $=(0,0).$
In other words, the matrix $U(s_{i})^{\epsilon_{i}}$ is cancelled at the $n_{i}$-th iteration. If it is cancelled by a matrix originally positioned in the first trace, then there exists $j\neq i$ such that either we have:
[…]
such that $f_{n_{i}}(\ell_{n_{i}-1}^{\prime}(j),j_{n_{i}-1}^{\prime}(j))=f_{n_{i}}(\ell_{n_{i}-1}(i),j_{n_{i}-1}(i))\pm 1$ and $U(s_{1,m+1-i})^{\epsilon_{1,m+1-i}}U(s_{1,m+1-j})^{\epsilon_{1,m+1-j}}=U(s_{i})^{\epsilon_{i}}U(s_{j})^{\epsilon_{j}}=\mathrm{Id}$. This is the case when the matrix originally at position $(1,m+1-i)$ is cancelled by a matrix originally at position $(1,m+1-j)$ in the same trace and with $\epsilon_{j}=-\epsilon_{i}$. In particular, we have $s_{i}=s_{j}$. Otherwise, and since we assume there is no rung cancellation of matrix $i$, there exists $j\neq i$ such that:
[…]
such that $f_{n_{i}}(\ell_{n_{i}-1}^{\prime}(j),j_{n_{i}-1}^{\prime}(j))=f_{n_{i}}(\ell_{n_{i}-1}(i),j_{n_{i}-1}(i))\pm 1$ and $U(s_{1,m+1-i})^{\epsilon_{1,2m+1-i}}U(s_{2,j})^{\epsilon_{2,j}}=U(s_{i})^{\epsilon_{i}}U(s_{j})^{-\epsilon_{j}}=\mathrm{Id}$. In particular once again $s_{i}=s_{j}$. The fact that all of the above remains true for $(s_{j}^{\prime},\epsilon_{j}^{\prime})_{j\in[m]}\sim_{\mathrm{P}}\mathcal{S}$ is a consequence of the previous construction of the function $f:[m]\rightarrow[m]$ that depended only on the movements of matrices and therefore the functions $f_{b}$ and $f_{b}^{rs}$ for $1\leq b\leq n$.
Remark 4.2.
In the previous proof we used, without mentioning it, the fact that for all $s\neq s^{\prime}$ for all $\epsilon,\epsilon^{\prime}\in\{+,-\}$ we have $U(s)^{\epsilon}U(s^{\prime})^{\epsilon^{\prime}}\neq\operatorname{Id}$ and $U(s)^{2}\neq\operatorname{Id}$ with probability one.
Lemma 4.9.
Let $\mathrm{P}$ a pattern of length $n$. Let $m\geq 1$ be an integer. We have for all $\mathcal{S}\in\mathcal{W}^{\prime}(m,\mathrm{P})$:
$|[\mathcal{S}]_{\mathrm{P}}|\leq d^{m/2}$
and:
$|\mathcal{W}^{\prime}(m,\mathrm{P})/\sim_{\mathrm{P}}|\leq(2m)^{2n}.$
Likewise let $p\geq 1$ be an integer. We have for all $\mathcal{W}^{\prime}_{1}(p,\mathrm{P})$:
$|[\mathcal{S}]_{\mathrm{P}}|\leq\frac{d}{d-1}(d-1)^{p}$
and:
$|\mathcal{W}_{1}^{\prime}(p,\mathrm{P})/\sim_{\mathrm{P}}|\leq(4p)^{2n}.$
Proof.
We consider $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$ a pattern and $m,p\geq 1$ integers fixed. We start by bounding the cardinal of $[\mathcal{S}]_{\mathrm{P}}$ in both cases. Let $\mathcal{S}\in\mathcal{W}^{\prime}(m,\mathrm{P})$. Considering Lemma 4.8, there exists $f:[m]\rightarrow[m]$ with no fixed point such that for all $1\leq i\leq[m]$ and for all $\mathcal{S}^{\prime}=(s_{1}^{\prime},\epsilon_{j}^{\prime})_{j\in[m]}\in[S]_{\mathrm{P}}$ we have $s_{f(i)}^{\prime}=s_{i}^{\prime}$. Since there is no fixed point it means that we have a subset $F\subset[m]$ that verifies $|F|\leq m/2$ and such that one only need to know $s_{j}^{\prime}$ for $j\in F$ to know all the values $s_{j}^{\prime}$ for $1\leq j\leq m$. Therefore we have $|[\mathcal{S}]_{\mathrm{P}}|\leq d^{m/2}$. For $\mathcal{S}\in\mathcal{W}_{1}^{\prime}(p,\mathrm{P})$ we will construct the subset $F$ as follows. Let $f:[2p]\rightarrow[2p]$ the function given by Lemma 4.8. We consider $F=F_{p}$ the set constructed as follows:
$F_{1}$ =$=\{1\}$
$\text{For all}\leavevmode\nobreak\ 2\leq t\leq p,\leavevmode\nobreak\ \leavevmode\nobreak\ F_{t}$ $=F_{t-1}\cup\{\inf\{i\notin\{F_{t-1};\leavevmode\nobreak\ \leavevmode\nobreak\ f(i)\notin f(F_{t-1})\}\}.$
We set $F=\{1<i_{2}<\cdots<i_{p^{\prime}}\}$ for $|F|=p^{\prime}\leq p$. We are left with counting the possible $s_{i_{t}}\in[d]$ for all $1\leq t\leq 2p$. Also in this case, by minimality we have that for all $1\leq t\leq 2p$, $s_{t+1}\neq s_{t}$. In particular we have necessarily $2\in F$. Therefore we have $d$ choices for $s_{1}$ and $d-1$ choices for $s_{2}$. Likewise for all $2\leq t\leq p^{\prime}$ we have at $d-1$ choices for $s_{i_{t}}$. Indeed $s_{i_{t}-1}$ is determined either by choice if $i_{t}-1\in F$, either by the function $f$ since there exists $i\in F_{t-1}$ such that $f(i)=i_{t}-1$ and $s_{i_{t}}\neq s_{i_{t}-1}$. It gives that $|[\mathcal{S}]_{\mathrm{P}}|\leq d(d-1)^{p^{\prime}-1}$, hence the result.
Now to upper-bound $|\mathcal{W}^{\prime}(m,\mathrm{P})/\sim_{\mathrm{P}}|$ we consider for all $1\leq b\leq n$:
$T_{b}(\mathrm{P})=\{(f_{t}(\mathcal{S},\mathrm{P}),f_{t}^{rs}(\mathcal{S},\mathrm{P}))_{t\in[b]};\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{S}\in\mathcal{W}^{\prime}(m)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ e(\mathcal{S},\mathrm{P})\neq\emptyset,e(\mathcal{S},\mathrm{P})\in\mathcal{R}_{n}\},$
where the tracking functions are defined by the algorithm if Section 3 applied to the word $\mathcal{S}$ in regard of the pattern $\mathrm{P}$. We consider the application:
$\phi_{b}:T_{b}$ $\rightarrow T_{b-1}\times[2m]\times[2m]$
$F=(F_{t})_{t\in[b]}$ $\mapsto((F_{t})_{t\in[b-1]},c_{1}(F),c_{2}(F))$
where we define $c_{1},c_{2}$ as the application that count the number of cancellations between generation $b-1$ and generation $b$ as follows. Let $F=(F_{t})_{t\in[b]}=(f_{t},f^{rs}_{t})_{t\in[b]}\in T_{b}$ and $\mathcal{S}\in\mathcal{W}^{\prime}(m,\mathrm{P})$ such that $F=(f_{t}(\mathcal{S},\mathrm{P}),f_{t}^{rs}(\mathcal{S},\mathrm{P}))_{t\in[b]}$. For all generations $1\leq b\leq n$, we use the notations introduced by Equation 26 in the following description. Then we have:
  • If $\ell_{b}=1$, $\epsilon_{b}=+$, then necessarily we have $\epsilon^{(b-1)}_{11}=\epsilon^{(b-1)}_{\ell_{b}j_{b}}$. It corresponds in the algorithm to the case where at the $b$-th iteration of Schwinger-Dyson equation we chose a term of the first line (17). Then necessarily we have $k_{b}=k_{b-1}+1$ and no possible cancellation of matrices in the following generation because of the hypothesis of minimality of words in the algorithm. Indeed we necessarily have $(s^{(b-1)}_{11},\epsilon_{11}^{(b-1)})\neq(s^{(b-1)}_{1j_{b}-1},-\epsilon_{1j_{b}-1}^{(b-1)})$ since we have $(s^{(b-1)}_{11},\epsilon_{11}^{(b-1)})=(s^{(b-1)}_{1j_{b}},\epsilon_{1j_{b}}^{(b-1)})$ and we suppose the $k_{b-1}$-word of generation $b-1$ minimal. We therefore set $c_{1}(F)=c_{2}(F)=0$.
  • If $\ell_{b}=1$, $\epsilon_{b}=-$ then we have $\epsilon^{(b-1)}_{11}=\epsilon^{(b-1)}_{\ell_{b}j_{b}}$ we still have $k_{b}=k_{b-1}+1$. It corresponds to the case where we chose the second line (18). Also now we have possible cancellations in the traces indexed $1$ and $2$. This number of cancellations is given by $c_{1}(F)=j_{b}-m^{(b)}_{1}=\min\{1\leq j\leq j_{b}-1;f_{b}^{rs}(1,j)=(0,0)\}$ and $c_{2}(F)=m_{1}^{(b-1)}-j_{b}-m_{2}^{(b)}=\max\{j_{b}+1\leq j\leq m_{1}^{(b-1)};f_{b}^{rs}(1,j)=(0,0)\}-j_{b}$, respectively the number of cancellations in the first and second traces.
  • $\ell_{b}\geq 2$, $\epsilon_{b}=+$ then we have $\epsilon_{11}^{(b-1)}=\epsilon^{(b-1)}_{\ell_{b}j_{b}}$ and $k_{b}=k_{b-1}-1$. It corresponds to the case where we chose the third line (19). As for the first case, the hypothesis of minimality implies that there are no cancellations possible. Therefore we set $c_{1}(F)=c_{2}(F)=0$.
  • Finally if $\ell_{b}\geq 2$ and $\epsilon_{b}=-\epsilon_{11}$ we have either $k^{\prime}=k-1$ or $k^{\prime}=k-2$. It corresponds to the case where we chose the fourth line (20). In this case we have cancellations in two different places in a same trace. We set in this case $c_{1}(F)=\min\{1\leq j\leq m_{1}^{(b-1)};\leavevmode\nobreak\ \leavevmode\nobreak\ f^{rs}_{b}(1,j)\neq(0,0)\}$ and $c_{2}(F)=\max\{1\leq j\leq m_{1}^{(b-1)};\leavevmode\nobreak\ \leavevmode\nobreak\ f^{rs}_{b}(1,j)\neq(0,0)\}$.
The function $\phi_{b}$ thus defined is injective. Therefore we have $|T_{b}|\leq(2m)^{2}|T_{b-1}|$ and therefore for all $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$ pattern we have:
$|\mathcal{W}^{\prime}(m,\mathrm{P})/\sim_{\mathrm{P}}|=|T_{n}|\leq(2m)^{2n}.$
The same argumentation gives the result for $\mathcal{W}_{1}^{\prime}(p,\mathrm{P})/\sim_{\mathrm{P}}$. ∎
5 Proof of Proposition 1.7
In all the following proof, for any $m\geq 1$ integer and any $(s_{j})_{j\in[2m]}\in[d]^{2m}$, we set $\mathcal{S}$ as the word given by:
$\mathcal{S}:=((s_{2m},+),(s_{2m-1},+),...,(s_{m+1},+),(s_{m},-),(s_{m-1},-),...,(s_{1},-)).$
and if we refer to $(s_{j})_{j\in[2m]}$ as a word we refer as the word obtained by considering the previous equation. For instance we have:
$1+\mathbb{E}(|\lambda_{2}|^{2m})\leq\left(\frac{1}{d}\right)^{2m}\sum_{\mathcal{S}\in[d]^{2m}}\mathbb{E}_{0}(\mathcal{S})=\mathbb{E}_{1}$
where for all $(s_{j})_{j\in[2m]}\in[d]^{2m}$, $\mathcal{S}$ is the word given right above. For all $1\leq m^{\prime}\leq m$ we denote by $\mathcal{W}_{sym}(m^{\prime})\subset\mathcal{W}(2m^{\prime})$ the set of $(s_{j})_{j\in[2m^{\prime}]}$ such that the corresponding minimal word is of length $2m^{\prime}$. Also we denote by $\mathcal{W}^{\prime}_{sym}(m^{\prime})\subset\mathcal{W}_{sym}(m^{\prime})\cap\mathcal{W}^{\prime}(2m^{\prime})$ the set of minimal word $\mathcal{S}\in\mathcal{W}_{sym}(m^{\prime})$ such that there exists $n\geq 1$ such that $\mathcal{R}_{n}(\mathcal{S})\neq\emptyset$, i.e. there exists a term in the algorithm that terminates after the $n$-th iteration with no rung cancellation and different from the empty set. We consider:
$m=\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor,$
and we re-write $\mathbb{E}_{1}$ as follows:
$\mathbb{E}_{1}$ $=\left(\frac{1}{d}\right)^{2m}\sum_{m^{\prime}=1}^{m}\sum_{\mathcal{S}\in\mathcal{W}_{sym}(m^{\prime})}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\begin{subarray}{c}(s^{\prime}_{j})_{j\in[2m]}\in[d]^{2m}\\ \mathcal{S}^{\prime}\sim\mathcal{S}\end{subarray}}\mathbb{E}_{0}(\mathcal{S})$
$=1+\sum_{m^{\prime}=1}^{m}d^{m-m^{\prime}}(m-m^{\prime}+1)\sum_{\mathcal{S}\in\mathcal{W}^{\prime}_{sym}(m^{\prime})}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{n=0}^{\infty}\sum_{e\in\mathcal{R}_{n}(\mathcal{S})}\bar{e}(\mathcal{S}).$
Between the first and second line we applied Lemma 4.6 and the fact for all $\mathcal{S}\in\mathcal{W}_{sym}(m^{\prime})$ we have:
$|(s_{j}^{\prime})_{j\in[2m]}\in[d]^{2m};\leavevmode\nobreak\ \mathcal{S}^{\prime}\sim\mathcal{S}\}|=\sum_{k=0}^{m-m^{\prime}}d^{k}d^{m-m^{\prime}-k}=d^{m-m^{\prime}}(m-m^{\prime}+1).$
Indeed it comes down to count the number of $(s_{j}^{\prime})_{j\in[2m]}\in[d]^{2m}$ such that we have:
$U(s_{2m}^{\prime})\cdots U(s_{m+1}^{\prime})U(s^{\prime}_{m})^{\ast}\cdots U(s_{1}^{\prime})^{\ast}\sim U(s_{2m^{\prime}})\cdots U(s_{m^{\prime}+1})U(s_{m^{\prime}})^{\ast}\cdots U(s_{1})^{\ast}$
where for $(t_{1},...,t_{k})\in[d]^{k}$ the equivalent class of $U(t_{1})\cdots U(t_{k})$ is $\{U(t_{1+o})U(t_{2+o})\cdots U(t_{k+o});\leavevmode\nobreak\ o\in\operatorname{\mathbb{N}}\}$. It means that for all such $(s_{j}^{\prime})_{j\in[2m]}$ we have:
$U(s_{1}^{\prime})^{\ast}\cdots U(s_{k}^{\prime})^{\ast}U(s_{2m-k+1}^{\prime})U(s_{2m-k+2}^{\prime})\cdots U(s_{2m}^{\prime})$ $=\operatorname{Id}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ U(s_{m+p}^{\prime})U(s_{m+p-1}^{\prime})\cdots U(s_{m+1}^{\prime})U(s_{m}^{\prime})^{\ast}U(s_{m-1})^{\ast}\cdots U(s_{m-p+1})^{\ast}$ $=\operatorname{Id}$
for some $(k,p)$ integers such that $p+k=m-m^{\prime}$. Therefore for $(k,p)$ fixed we have $d^{k}d^{p}$ possible $(s_{j}^{\prime})_{j\in[2m]}$. We now re-write $\mathbb{E}_{1}$:
$\mathbb{E}_{1}$ $\leq 1+\left(\frac{1}{d}\right)^{2m}m\sum_{m^{\prime}=1}^{m}\sum_{\mathcal{S}\in\mathcal{W}_{sym}^{\prime}(m^{\prime})}d^{m-m^{\prime}}\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\mathds{1}(e(\mathcal{S},\mathrm{P})\neq\emptyset)N^{2-2/3n}$
where the last summand is over all patterns $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}\in([4m]\times[4m]\times\{+,-\})^{n}$ and where we applied Lemma 3.7 to bound $|\bar{e}(\mathcal{S},\mathrm{P})|$ for $e(\mathcal{S},\mathrm{P})\in\mathcal{R}_{n}(\mathcal{S})$. Exchanging the summand in the previous equation we have:
$\mathbb{E}_{1}$ $\leq 1+\left(\frac{1}{d}\right)^{2m}m\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\sum_{m^{\prime}=1}^{m}d^{m-m^{\prime}}\sum_{[\mathcal{S}]_{\mathrm{P}}\in\mathcal{W}^{\prime}(2m^{\prime},\mathrm{P})/\sim_{\mathrm{P}}}|[\mathcal{S}]_{\mathrm{P}}|N^{2-2/3n}.$
Applying Lemma 4.9 we have:
$\mathbb{E}_{1}$ $\leq 1+\left(\frac{1}{d}\right)^{2m}m\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\sum_{m^{\prime}=1}^{m}d^{m-m^{\prime}}(4m)^{2n}d^{m^{\prime}}N^{2-2/3n}$
$\leq 1+\left(\frac{1}{d}\right)^{2m}m^{2}N^{2}d^{m}\sum_{n=0}^{\infty}\left(\frac{512m^{4}}{N^{1/3}}\right)^{n}$
$\leq 1+\rho_{d}^{2m}2m^{2}N^{2},$
where the last inequality is obtained by replacing $m$ by its value, that is:
$m=\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor.$
We consider $N$ large enough so we have $m\geq\frac{1}{8\sqrt{2}}N^{1/12}$. Let now $\epsilon>0$, replacing $m$ by its expression we have that:
$\frac{\mathbb{E}(s_{2}(\mathcal{E}^{m})^{2})}{\rho_{d}^{2m}(1+\epsilon)^{2m}}$ $\leq\frac{\mathbb{E}_{1}-1}{(1+\epsilon)^{2m}\rho_{d}^{2m}}$
$\leq\frac{2m^{2}N^{2}}{(1+\epsilon)^{2m}}\leq\frac{N^{13/6}}{8}e^{-\frac{1}{4\sqrt{2}}\ln(1+\epsilon)N^{1/12}}$
which implemented in (5) ends the proof of Proposition 1.7.
6 Proof of Theorem 1.8
In all the following proof, for any $m,p\geq 1$ integers and any $(s_{j}^{t})_{j\in[m],t\in[2p]}\in[d]^{2mp}$, we set $\mathcal{S}$ as the word given by:
$\mathcal{S}:=((s_{m}^{2p},+),(s_{m-1}^{2p},+),...,(s_{1}^{2p},+),(s_{m}^{2p-1},-),(s_{m-1}^{2p-1},-),...,(s_{j}^{t},\epsilon_{t}),(s_{j-1}^{t},\epsilon_{t}),...,(s_{1}^{1},-))$
where for all $1\leq t\leq 2p$ we set $\epsilon_{t}=+$ if $t$ is even and $\epsilon_{t}=-$ if $t$ is odd. Again if we refer to $(s_{j}^{t})_{j\in[m],t\in[2p]}$ as a word we refer as the word obtained by considering the previous equation. For instance we have:
$1+\mathbb{E}(\lambda_{2}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p}))\leq\mathbb{E}\left(\sum_{a=1}^{N^{2}}\lambda_{a}(({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p})\right)=\left(\frac{1}{d}\right)^{2mp}\sum_{(s^{t}_{j})\in[d]^{2mp}}\mathbb{E}_{0}(\mathcal{S})=\mathbb{E}_{2}.$
We apply Proposition 3.6 and Lemma 4.6 and we re-write $\mathbb{E}_{2}$:
$\mathbb{E}_{2}$ $=\left(\frac{1}{d}\right)^{2mp}\sum_{m^{\prime}=1}^{mp}\sum_{\mathcal{S}\in\mathcal{W}(m^{\prime})}\sum_{\begin{subarray}{c}\mathcal{S}^{\prime}\in[d]^{2mp}\\ \mathcal{S}^{\prime}\sim\mathcal{S}\end{subarray}}\mathbb{E}_{0}(\mathcal{S})$
$=1+\sum_{m^{\prime}=1}^{mp}\sum_{\mathcal{S}\in\mathcal{W}^{\prime}(m^{\prime})}|\underbrace{\{\mathcal{S}^{\prime}\in[d]^{2mp};\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{S}^{\prime}\sim\mathcal{S}\}}_{\mathcal{A}(m,p,\mathcal{S})}|\sum_{n=0}^{\infty}\sum_{e\in\mathcal{R}_{n}(\mathcal{S})}\bar{e}(\mathcal{S}).$ (29)
We will now upper bound the cardinal of $\mathcal{A}(m,p,\mathcal{S})$ independently of the choice $\mathcal{S}$ minimal word. We first fix $(m_{1},...,m_{2p})\in\{0,...,m\}^{2p}$ such that $\sum_{t}m_{t}=2m^{\prime}$ and for all $1\leq t\leq m_{2p}$, there exists $(\tilde{s}_{j}^{t})\in[d]^{m_{t}}$ such that:
$\mathcal{U}(\mathcal{S})=U(\tilde{s}_{m_{1}}^{1})\cdots U(\tilde{s}_{1}^{1})U(\tilde{s}_{m_{2}}^{2})^{\ast}\cdots U(\tilde{s}_{1}^{2})^{\ast}\cdots U(\tilde{s}_{m_{2p-1}}^{2p-1})U(\tilde{s}^{2p}_{m_{2p}})^{\ast}\cdots U(\tilde{s}^{2p}_{1})^{\ast}.$
where we take as convention that the product over an empty set is $\prod_{\emptyset}=\operatorname{Id}$. For all minimal $\mathcal{S}\in\mathcal{W}(m^{\prime})$ we have at most $(m+1)^{2p}$ ways to chose $(m_{t})_{t\in[2p]}$. For such a choice of $(m_{t})_{t}$ we consider a choice of $0\leq k_{1}\leq m-m_{1}$ from which we define the sequence:
$k_{t+1}=m-m_{t}-k_{t}.$
Once again for $\mathcal{S}$ minimal and $(m_{t})_{t}$ fixed we have at most $(m+1)$ choices. Finally we consider all the words $\mathcal{S}^{\prime}=(s_{j}^{t})_{j\in[m],t\in[2p]}$ such that for all $1\leq t\leq 2p$ and for all $1\leq j\leq m_{t}$ we have:
$(s_{k_{t}+j}^{t},\epsilon_{k_{t}+j}^{t})=(\tilde{s}^{t}_{j},\epsilon_{t}).$
In other words for all $2\leq t\leq 2p$ we have:
$\cdots U(s^{t-1}_{1})^{\epsilon_{t-1}}\cdots\underbrace{U(s^{t-1}_{k_{t-1}+1})^{\epsilon_{t}}\cdots U(s^{t-1}_{k_{t-1}+m_{t-1}})^{\epsilon_{t-1}}}_{=U(\tilde{s}_{1}^{t-1})^{\epsilon_{t-1}}\cdots U(\tilde{s}_{m_{t-1}}^{t-1})^{\epsilon_{t-1}}}\underbrace{U(s^{t-1}_{k_{t-1}+m_{t-1}+1})^{\epsilon_{t-1}}\cdots U(s_{m}^{t-1})^{\epsilon_{t-1}}U(s_{1}^{t})^{\epsilon_{t}}\cdots U(s_{k_{t}}^{t})^{\epsilon_{t}}}_{=\operatorname{Id}}\cdots.$
For $\mathcal{S}$ minimal, $(m_{t})_{t\in[2p]}$ and $(k_{t})_{t\in[2p]}$ fixed we can construct at most $\prod_{t=1}^{2p}d^{k_{t}}=d^{mp-m^{\prime}}$ corresponding equivalent word. We may now re-write $\mathbb{E}_{2}$ as follows
$\mathbb{E}_{2}$ $\leq 1+\left(\frac{1}{d}\right)^{2mp}\sum_{m^{\prime}=1}^{mp}\sum_{\mathcal{S}\in\mathcal{W}^{\prime}(m^{\prime})}(m+1)^{2p+1}d^{mp-m^{\prime}}\sum_{n=0}^{\infty}\sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\mathds{1}(e(\mathcal{S},\mathrm{P})\neq\emptyset)N^{2-2/3n}.$
Exchanging the summand we have:
$\mathbb{E}_{2}$ $\leq 1+\left(\frac{1}{d}\right)^{2mp}\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\sum_{m^{\prime}=1}^{mp}(m+1)^{2p+1}d^{mp-m^{\prime}}\sum_{[\mathcal{S}]_{\mathrm{P}}\in\mathcal{W}^{\prime}(m^{\prime},\mathrm{P})/\sim(\mathrm{P})}|[\mathcal{S}]_{\mathrm{P}}|N^{2-2/3n}$
Applying Lemma 4.9 we have that for all pattern $\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}$:
$|[\mathcal{S}]_{\mathrm{P}}|\leq d^{m^{\prime}}$
$|\mathcal{W}^{\prime}(m^{\prime},\mathrm{P})/\sim(\mathrm{P})|\leq(4mp)^{2n}.$
Therefore we have:
$\mathbb{E}_{2}$ $\leq 1+\left(\frac{1}{d}\right)^{2mp}\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \left(16(mp)^{2}\right)^{n}\sum_{m^{\prime}=1}^{mp}(m+1)^{2p+1}d^{mp-m^{\prime}}(4mp)^{2n}d^{m^{\prime}}N^{2-2/3n}$
$\leq 1+4(m+1)^{2p+1}pN^{2}\rho_{d}^{2mp}\sum_{n=0}^{\infty}\left(\frac{512(mp)^{4}}{N^{1/3}}\right)^{n}$
We now consider $1\leq m\leq\frac{1}{2}\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$. We set:
$p=\lfloor\frac{1}{m}\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor\rfloor\leq\frac{1}{m}\frac{1}{4\sqrt{2}}N^{1/12}$
which implies that $\frac{512(mp)^{4}}{N^{1/3}}\leq 1/2$. Therefore we have:
$\frac{\mathbb{E}(\lambda_{2}({\mathcal{E}^{\ast}}^{m}\mathcal{E}^{m})^{p})}{\rho_{d}^{2mp}(1+\epsilon)^{2mp}}\leq\frac{\mathbb{E}_{2}-1}{\rho_{d}^{2mp}(1+\epsilon)^{2mp}}\leq\frac{8(m+1)^{2p+1}pN^{2}}{(1+\epsilon)^{2mp}}.$
For $N$ large enough, for all sequence $m\leq\frac{1}{2}\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$ and $p=p(m,N)$ defined above we have:
$mp\geq\frac{1}{8\sqrt{2}}N^{1/12}.$
Therefore for all $\epsilon>0$, for all $N$ and all $1\leq m\leq\frac{1}{2}\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$, replacing $p$ by its value we have:
$\frac{8(m+1)^{2p+1}pN^{2}}{(1+\epsilon)^{2mp}}\leq\frac{\sqrt{2}(m+1)}{m}N^{25/12}e^{[2\frac{\ln(m+1)}{m}-\ln(1+\epsilon)]\frac{1}{4\sqrt{2}}N^{1/12}}.$
which implemented in Equation 9 concludes the proof.
7 Proof of Theorem 1.9
To prove the last Theorem we need to look further in the combinatorial computations of the previous proof. Indeed we consider Equation (29) for $m=1$, that is:
$\mathbb{E}_{2}=1+\left(\frac{1}{d}\right)^{2p}\sum_{p^{\prime}=1}^{p}\sum_{\mathcal{S}\in\mathcal{W}^{\prime}(p^{\prime})}|\underbrace{\{\mathcal{S}^{\prime}\in[d]^{2p};\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{S}^{\prime}\sim\mathcal{S}\}}_{\mathcal{A}(p,\mathcal{S})}|\sum_{n=0}^{\infty}\sum_{e\in\mathcal{R}_{n}(\mathcal{S})}\bar{e}(\mathcal{S}).$ (30)
We upper bound $|\mathcal{A}(p,\mathcal{S})|$ independently of the choice of $\mathcal{S}$ minimal. Therefore we may suppose that $\mathcal{S}$ starts with $\epsilon_{1}=+$, i.e. we can write $((\tilde{s}_{1},+),(\tilde{s}_{2},-),...,(\tilde{s}_{2p^{\prime}},-))=\mathcal{S}$. First notice that we have:
$|\mathcal{A}(p,\mathcal{S})|\leq 2p|\{\mathcal{S}=(s_{1},...,s_{2p});\leavevmode\nobreak\ \leavevmode\nobreak\ U(s_{1})U(s_{2})^{\ast}\cdots U(s_{2p-1})U(s_{2p})^{\ast}=\mathcal{U}(\mathcal{S})\}|=:|\mathcal{A}^{\prime}(p,\mathcal{S})|.$
Indeed if $U(s_{1})U(s_{2})^{\ast}\cdots U(s_{2p-1})U(s_{2p})^{\ast}=\mathcal{U}(\mathcal{S})$ then for all $1\leq t\leq 2p$ we have $(s_{t+1},s_{t+2},...,s_{2p+t})\sim\mathcal{S}$ where we recall that by convention $s_{2p+1}=s_{1}$. Now to upper bound $|\mathcal{A}^{\prime}(p,\mathcal{S})|$ we consider the following random walk:
$(s_{t})_{t\in\mathbb{N}}$ $\sim\mathcal{U}([d])$
$X_{0}$ $:=\operatorname{Id}$
$\forall t\in\operatorname{\mathbb{N}},\leavevmode\nobreak\ \forall s\in[d]\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{P}(X_{2t+1}=X_{2t}U(s))=\frac{1}{d},\leavevmode\nobreak$ $\leavevmode\nobreak\ \mathbb{P}(X_{2t+2}=X_{2t+1}U(s)^{\ast})=\frac{1}{d}.$
[…]
$\mathbb{P}(\ell_{2p}=2p^{\prime})=\frac{N(p,p^{\prime})}{d^{2p}}\leq\frac{2^{2p}}{d^{2p}}(d-1)^{p+p^{\prime}}.$
Also for all $t$ integer we denote by $\mathrm{Red}(2p^{\prime})=d(d-1)^{2p^{\prime}-1}$ the number of reduced words of length $2p^{\prime}$ that the previous random walk can have, i.e. the number of $(s_{1}^{\prime},...,s_{2p^{\prime}}^{\prime})\in[d]^{2p^{\prime}}$ such that for all $1\leq t\leq 2p^{\prime}$ we have $s_{t}^{\prime}\neq s_{t+1}^{\prime}$ and we have:
$N(p,p^{\prime})=|\mathcal{A}^{\prime}(p,\mathcal{S})|\mathrm{Red}(2p^{\prime})$
and therefore:
$|\mathcal{A}(p,\mathcal{S})|\leq 2p2^{2p}\frac{d-1}{d}(d-1)^{p-p^{\prime}}.$
$\mathbb{E}_{3}$ $\leq 1+\left(\frac{1}{d}\right)^{2p}\sum_{p^{\prime}=1}^{p}\sum_{\mathcal{S}\in\mathcal{W}_{1}^{\prime}(p^{\prime})}2p2^{2p}\frac{d-1}{d}(d-1)^{p-p^{\prime}}\sum_{n=0}^{\infty}\sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\mathds{1}(e(\mathcal{S},\mathrm{P})\neq\emptyset)N^{2-2/3n}.$
Exchanging the summand we have:
$\mathbb{E}_{3}$ $\leq 1+\left(\frac{1}{d}\right)^{2p}\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{\mathrm{P}=(\ell_{b},j_{b},\epsilon_{b})_{b\in[n]}}\sum_{p^{\prime}=1}^{p}2p2^{2p}\frac{d-1}{d}(d-1)^{p-p^{\prime}}\sum_{[\mathcal{S}]_{\mathrm{P}}\in\mathcal{W}_{1}^{\prime}(p^{\prime},\mathrm{P})/\sim(\mathrm{P})}|[\mathcal{S}]_{\mathrm{P}}|N^{2-2/3n}.$
Applying Lemma 4.9 we have:
$\mathbb{E}_{3}$ $\leq 1+\left(\frac{1}{d}\right)^{2p}\sum_{n=0}^{\infty}\leavevmode\nobreak\ \leavevmode\nobreak\ \left((4p)^{2}2\right)^{n}\sum_{m^{\prime}=1}^{mp}2p2^{2p}\frac{d-1}{d}(d-1)^{p-p^{\prime}}(4p)^{2n}(d-1)^{p^{\prime}}\frac{d}{d-1}N^{2-2/3n}$
$\leq 1+2pN^{2}\sigma_{d}^{2p}\sum_{n=0}^{\infty}\left(\frac{512p^{4}}{N^{1/3}}\right)^{n}.$
We consider:
$p:=\lfloor\frac{1}{4\sqrt{2}}N^{1/12}\rfloor$
and obtain:
$\mathbb{E}_{3}\leq 1+4pN^{2}\sigma_{d}^{2p}.$
Finally we have:
$\frac{\mathbb{E}(\lambda_{2}({\mathcal{E}^{\ast}}\mathcal{E})^{p})}{\sigma_{d}^{2p}(1+\epsilon)^{2p}}\leq\frac{\mathbb{E}_{3}-1}{\sigma_{d}^{2p}(1+\epsilon)^{2p}}\leq\frac{4pN^{2}}{(1+\epsilon)^{2p}}.$